### v1

# The JADES Origins Field: A New JWST Deep Field in the JADES Second NIRCam Data Release

Abstract
We summarize the properties and initial data release of the JADES Origins Field (JOF), which will soon be the deepest imaging field yet observed with the James Webb Space Telescope (JWST). This field falls within the GOODS-S region about 8’ south-west of the Hubble Ultra Deep Field (HUDF), where it was formed initially in Cycle 1 as a parallel field of HUDF spectroscopic observations within the JWST Advanced Deep Extragalactic Survey (JADES). This imaging will be greatly extended in Cycle 2 program 3215, which will observe the JOF for 5 days in six medium-band filters, seeking robust candidates for galaxies. This program will also include ultra-deep parallel NIRSpec spectroscopy (up to 104 hours on-source, summing over the dispersion modes) on the HUDF. Cycle 3 observations from program 4540 will add 20 hours of NIRCam slitless spectroscopy to the JOF. With these three campaigns, the JOF will be observed for 380 open-shutter hours with NIRCam using 15 imaging filters and 2 grism bandpasses. Further, parts of the JOF have deep 43 hr MIRI observations in F770W. Taken together, the JOF will soon be one of the most compelling deep fields available with JWST and a powerful window into the early Universe. This paper presents the second data release from JADES, featuring the imaging and catalogs from the year 1 JOF observations.
1 Introduction
A primary goal of the James Webb Space Telescope (JWST) is to study galaxies to the earliest observable epochs, when the first massive assemblies of gas and stars set in motion the long play of galaxy evolution. Precisely timing when the first galaxies formed after the Big Bang is a key goal of observational cosmology and one of the last great unknowns in our history of the Universe: When and how did the first galaxies form? The primordial star formation and the formation of galaxies in the first 300 Myr (redshift ) after the Big Bang is a sensitive probe of the physics governing the conversion of gas into stars at early times and different dark matter models (e.g., Behroozi & Silk, 2015; Dayal & Ferrara, 2018; Boylan-Kolchin, 2023; Kannan et al., 2023). Additionally, these first, unexplored galaxies might harbor surprises such as metal-free stellar populations (Population III) or rapidly accreting young black hole seeds, thereby shedding light onto possible pathways to the formation of the billion-solar-mass black holes observed at (e.g., Bromm & Yoshida, 2011; Xu et al., 2013; Smith & Bromm, 2019; Vikaeus et al., 2022; Volonteri et al., 2021; Trinca et al., 2023; Schneider et al., 2023). Deep fields, both in imaging and associated spectroscopy, are a key opportunity for this work, allowing the astronomy community to coordinate its multi-wavelength resources to best advantage.
Cycle 1 observations with JWST succeeded quickly in pushing the redshift frontier beyond what had been seen with HST. Relying on the complete scattering of photons shortward of Ly 1216Å by the intergalactic medium (IGM), many F115W dropouts at and even some F150W dropouts at were detected as photometric candidates in numerous fields (e.g., Bouwens et al., 2022; Castellano et al., 2022; Finkelstein et al., 2022; Harikane et al., 2023a; Naidu et al., 2022b; Adams et al., 2023; Austin et al., 2023; Atek et al., 2023; Donnan et al., 2023; Franco et al., 2023; Hainline et al., 2023; Wang et al., 2023a; Whitler et al., 2023; Yan et al., 2023). As of this writing, numerous galaxies at , including two at , have been spectroscopically confirmed with long NIRSpec exposures (Curtis-Lake et al., 2023; Heintz et al., 2023; Morishita & Stiavelli, 2023; Roberts-Borsani et al., 2023; Tang et al., 2023; Arrabal Haro et al., 2023b; Bunker et al., 2023b; Hsiao et al., 2023; Sanders et al., 2023b; Wang et al., 2023b; Fujimoto et al., 2023b), along with a larger number of galaxies out to where the much brighter [OIII] and H lines are available in the NIRSpec bands.
These early results have demonstrated not only JWST’s unrivalled capabilities in successfully exploring the early Universe, but also its potential for pushing the redshift frontier even further. Indeed, it is evident that JWST—with its excellent design and on-orbit performance (Rigby et al., 2022)—is fully capable of probing back to even earlier cosmic times. Within this context, we note that the Cycle 1 programs were designed to go deep in F115W and F150W, not in F200W. The few hints of any F200W dropouts are uncertain, potentially mimicked by –5 galaxies with dust-reddened continua and well-placed strong rest-optical emission lines (Fujimoto et al., 2023a; Naidu et al., 2022a; Zavala et al., 2022; Arrabal Haro et al., 2023a).
We describe here the construction of a new deep field, the JADES Origins Field (JOF), that aims to push our exploration of early galaxies further than ever before possible. This field was initiated by the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein et al., 2023) as JWST NIRCam coordinated parallel observations to a NIRSpec multi-object spectroscopy (MOS) pointing focused on the Hubble Ultra Deep Field (HUDF; Beckwith et al., 2006) within the GOODS-S field (Giavalisco et al., 2004). These Cycle 1 observations of JADES were described in Eisenstein et al. (2023) and Bunker et al. (2023a). New data in Cycle 2 and 3 will bring this field to nearly 10 mission days of NIRCam imaging and slitless spectroscopy in addition to over 7 mission days of deep MIRI imaging, making it the deepest JWST imaging yet achieved. The geometry of this field is well suited for future JWST work, as the parallels can span between the JADES Deep mosaic around the HUDF and the JOF at observing windows that have low zodiacal background. As a start on this, NIRSpec MOS work in the JOF will be conducted as part of JADES in late 2023.
In this paper, we detail the scientific design and implementation of the Cycles 2 and 3 observations. Relative to other deep fields with JWST, this combined data set features very deep medium-band imaging in a total of 15 NIRCam filters. The medium-band imaging is motivated by the goal of having finer sampling of the spectral energy distribution for searches for Ly dropouts at . As we argue in § 3 (and as mentioned above), broad-band NIRCam data alone is susceptible to selection of Balmer-break and emission-line interlopers from mid-redshift, and this failure mode appears to become more pernicious at , because of the rarity of these galaxies and because of the more limited number of applicable filters. Using medium-bands F162M, F182M, F210M, F250M, F300M, and F335M will exclude mid-redshift emission lines and allow for detection of more distinctive breaks. We expect that these data will provide robust observational constraints on the number density of galaxies and their rest-UV luminosities, colors, and sizes.
Of course, these exceptionally deep images, up to 46 hours per filter and covering about 9 square arcminutes, will be of high value for the study of intermediate-redshift galaxies as well. To support this, the NIRCam instrument team will conduct NIRCam slitless spectroscopy in Cycle 3 in both the F356W and F444W filters, reaching up to 9 hr of on-source integration time. These will provide spectroscopic redshifts for hundreds of galaxies at with the strong H and/or [OIII] 5007Å lines.
A further opportunity of the Cycle 1 and 2 data is the observation of a very deep NIRSpec MOS pointing on the UDF. The Cycle 1 data was already released in Bunker et al. (2023a); here we describe the science drivers and layout of the Cycle 2 pointing, which will target faint high-redshift galaxies with exposure times up to 104 hrs on-source.
In addition to an overview of the upcoming programs, this paper presents the initial release of NIRCam imaging data and catalogs on this field, already some of the deepest imaging obtained by JWST. The release includes medium-depth flanking fields that connect this footprint to the HUDF region released in the JADES first data release (Rieke et al., 2023).
In § 2–4, we present the Cycle 1, 2, and 3 programs in the JOF. § 5 then presents the initial data release, and we summarize in § 6.
2 Cycle 1: Beginning this field as a Coordinated Parallel
The imaging in this field began as the NIRCam Coordinated Parallel of program 1210 (PI: Lützgendorf), which executed the JADES Deep/HST NIRSpec micro-shutter array (MSA) observations in and around the HUDF (Bunker et al., 2023a). Fortuitously, the position angle of these observations put these NIRCam parallels on top of further data from program 1180 (PI: Eisenstein) that included medium-depth NIRCam imaging and deep MIRI F770W imaging, the latter resulting from coordinated parallels of deep NIRCam imaging around the HUDF. These observations are described in Eisenstein et al. (2023). In brief, the 1210 parallel provides 55 open-shutter dual-filter hours of NIRCam imaging in 9 filters: F090W, F115W, F150W, F200W, F277W, F335M, F356W, F410M, and F444W. The 1180 medium imaging overlaps and extends this geometry at flanking depth in all filters save F335M. The MIRI imaging will reach 43 hr depth in F770W in each of 4 adjoining fields.
Further, an assigned observing window in Year 2 has allowed us to place one Coordinated Parallel NIRCam field from the JADES Medium/JWST tier (program 1286, PI: Lützgendorf) in a partially overlapping location about 1 south of the 1210 pointing. This is 11 hours open-shutter of additional data in all the previous nine filters plus F070W. It will add further depth to the JOF imaging, but also will make a deeper region (being the superposition of two medium surveys) that is more square and hence better suited for the NIRSpec MOS footprint. The extra NIRCam imaging also overlaps more of the deep MIRI data.
A simplified view of the GOODS-S region is provided in Figure 1, showing the location of the JOF relative to the HUDF, MUSE, and other JWST programs.
3 Cycle 2: Deep Medium-Band Imaging with Program 3215
3.1 Science Motivation: Key Opportunities at
When, how, and why the first stars and black holes formed after the Big Bang is one of the provocative mysteries of the Universe. The galaxies hosting those stars and black holes have remained out of reach until now. With the advent of JWST, we now have a telescope in space that can directly probe and characterize the formation of those first galaxies. This is a key goal of JWST and observational cosmology because the first galaxies probe unknown physics related to star formation (e.g., the efficiency of how gas is converted into stars in early dark matter halos e.g., Behroozi & Silk 2015; Gnedin 2016; Vogelsberger et al. 2020; Kannan et al. 2023; Wilkins et al. 2023) and early structure formation (e.g., Dayal & Ferrara, 2018; Khimey et al., 2021; Lovell et al., 2023; Boylan-Kolchin, 2023). While the low metallicity in this first generation of galaxies should slow H formation and decrease star-formation efficiency (e.g., Krumholz & Dekel, 2012), the gas-rich, compact nature of early galaxies may result in increased gas surface densities and star-formation efficiencies (e.g., Krumholz et al., 2009; Ostriker & Shetty, 2011; Somerville et al., 2015).
There are a wide variety of theoretical predictions for the redshift evolution of the UV luminosity function of galaxies (e.g., Behroozi & Silk, 2015; Mason et al., 2015; Tacchella et al., 2018; Kannan et al., 2023; Wilkins et al., 2023; Katz et al., 2021) These different models typically lead to similar outcomes at regarding number counts and the cosmic star-formation rate density, but they diverge more and more toward higher redshifts due to different treatments of cooling, star formation efficiency, and feedback. Similarly, predictions for the formation and evolution of supermassive black holes vary widely (Habouzit et al., 2022). With JWST, we have the opportunity to push to the redshift frontier of where the physics of these extreme settings cause large differences, factors of 10 and more.
Of course, the redshift frontier offers the opportunity for surprises and more exotic discoveries. As dark matter halos are rapidly growing in this early epoch, galaxies at are a sensitive probe of the interplay of gravitational collapse, gas cooling, and energetic feedback. Indeed, this epoch could even reveal surprises in structure formation and the nature of dark matter: the suppression of small-scale structure in some dark matter models leads to a delayed and subsequently more rapid stellar assembly relative to cold dark matter models (e.g., Gandolfi+ 2022). The exploration of is the most direct way to constrain the onset of cosmic dawn. Together with the constraints on the star-formation efficiency and feedback processes at , JWST can bridge the gap to ongoing 21cm experiments—such as EDGES (Bowman et al., 2018), SARAS 3 (Singh et al., 2022) and REACH (de Lera Acedo et al., 2022)—and thereby creating new synergies to constrain the formation of the first stars and galaxies.
In addition to delivering number density constraints at , JWST observations will constrain the properties of these first galaxies. Specifically, these data are sensitive to the total rest-frame UV luminosity, UV continuum slope , and the morphology of those first structures. The UV continuum slope is a well-known diagnostic for the stellar populations in galaxies (Bouwens et al., 2009; Wilkins et al., 2011; Finkelstein et al., 2012; Bhatawdekar & Conselice, 2021), primarily affected by the dust attenuation and metallicity. Together with the UV luminosity, one can derive dust-corrected star-formation rates (SFRs) and infer the cosmic SFR density at . Furthermore, an estimate of will probe the chemical enrichment and dust physics in the earliest galaxies, extending studies at from lower redshifts (e.g., Ferrara et al., 2022). Very blue UV slopes are a potential signposts of Population III stars or accreting black holes (Schaerer, 2003; Dayal & Ferrara, 2012).
The structure and morphology of galaxies is a well-known tracer of galaxy assembly (e.g., Conselice, 2014). JWST images reveal marginally-resolved galaxies at with compact ( pc) sizes (Robertson et al., 2023; Ono et al., 2023). Deep imaging at will deliver rest-UV size constraints down to a resolution of 300 pc. Measurements of the size and shape (even upper limits) are a sensitive probe of stellar and black hole feedback in those galaxies (e.g., Pillepich et al., 2019; Wu et al., 2020) and could potentially probe the filamentary structure of the gas accretion (e.g., Pandya et al., 2019; Ceverino et al., 2023). Further, one will be sensitive to the multiplicity and close-pair fraction of galaxies at , probing the importance of mergers and star-formation in filaments in these early galaxies (e.g., Wang & Loeb, 2018; O’Leary et al., 2021).
3.2 Observational Challenges and a Method to Reach
There are important reasons why pushing for F200W () dropouts is substantially harder than F115W ( or F150W () dropouts: 1) Such galaxies are fainter and rarer. 2) The frequency of extreme emission-line galaxies increases with redshift, and JWST clearly has shown a substantial population of them at , where they can create SEDs that differ from those of purely stellar populations, boosting the rest-optical and enhancing the apparent Balmer drop. 3) The dichroic gap in NIRCam leaves a wider spacing between F200W and F277W that makes it harder to exclude a mildly reddened Balmer-drop galaxy at , particularly with [OIII] potentially enhancing F277W. 4) Many Cycle 1 programs include only F277W, F356W, F410M, and F444W.111Notable exceptions are JEMS (Williams et al., 2023) and CANUCS (Withers et al., 2023). However, the filters beyond 4 microns have reduced sensitivity due to increased zodiacal foreground, so faint galaxies may only be well-detected in the bluer two bands, which in turn can be filled with [OIII] and H. These higher-redshift break candidates are less compelling than lower redshift examples that have 3 or 4 sensitive bands redward of the proposed break.
Whereas Cycle 1 programs pushed deep in F115W and/or F150W, in program 3215 we adopt a strategy optimized to identify candidates at . In particular, this search must 1) push very deep in the 2 micron region; 2) increase the spectral resolution to isolate the dropout; and 3) include additional filters in the 2.5–4 micron region to eliminate contamination from mid-redshift emission-line galaxies. We believe that this combination of high-SNR detections, well-sampled dropout features, and systematic control over contaminants can reach the high standard of proof that claims of galaxies will require. The new imaging will achieve these requirements by combining the usual NIRCam wide-band filters with well-chosen medium-band filters. By observing very deep in F182M, F210M, and F250M, the photometry can isolate a sharp break within a narrower and hence more distinctive region. Further, using the combination of medium and wide-bands in the long-wave (LW) channels will reveal strong emission lines that could supply false positive Balmer breaks.
We illustrate this strategy by considering the candidate in the CEERS field (Donnan et al., 2023), scaled down by a factor of 30 in flux and assuming the noise levels of programs 1210 & 3215. While now proved to be at (Arrabal Haro et al., 2023a), this object was a plausible candidate, showing a drop in F200W by a factor of 4 relative to F277W, no detection in F150W, and a blue continuum in the LW filters. However, this photometry was also a match to models at with strong optical emission lines and reddening suppression in the UV (Naidu et al., 2022a). Figure 2 shows how these two hypotheses are starkly differentiated by the NIRCam medium bands: the Lyman break is much steeper than a dust-enhanced Balmer break, and the medium bands easily distinguish continuum from emission lines. Shifting the true galaxy in redshift and repeating the fits, Figure 3 demonstrates that this suite of filters provides very effective recovery of candidates at between 13 and 20.
3.3 Filters, Exposure Times, and Depths
| Filter | Cycle 2, Program 3215 | Cycle 1 Time | Total Time | 5- PS | Ly | |
| Time (ks) | (ks) | (ks) | (nJy) | redshift | ||
| F090W | — | — | 55.2 | 55.2 | 3.01 | |
| F115W | — | — | 72.9 | 72.9 | 2.38 | |
| F150W | — | — | 55.2 | 55.2 | 2.14 | |
| F162M | 30 | 82.5 | — | 82.5 | 2.62 | |
| F182M | 60 | 165.0 | — | 165.0 | 1.59 | |
| F200W | — | — | 38.7 | 38.7 | 2.25 | |
| F210M | 45 | 123.8 | — | 123.8 | 2.15 | |
| F250M | 60 | 165.0 | — | 165.0 | 2.40 | |
| F277W | — | — | 47.0 | 47.0 | 1.90 | |
| F300M | 45 | 123.8 | — | 123.8 | 1.80 | |
| F335M | 30 | 82.5 | 24.8 | 107.3 | 1.81 | |
| F356W | — | — | 38.7 | 38.7 | 1.95 | |
| F410M | — | — | 55.2 | 55.2 | 3.30 | |
| F444W | — | — | 56.4 | 56.4 | 2.65 | |
Because of the opportunity to re-use the deep imaging of program 1210, we designed program 3215 to return to the same footprint, matching the position angle exactly. The imaging exposure times and depths, in both new and old filters, are listed in Table 1, along with the Ly redshift range for each.
The driving requirement is to go deep in F182M and F210M, as potential dropout bands. In total, F182M, F200W, and F210M will reach 91 hrs of exposure time, far more than the 11 hrs that 1210 & 1180 provide and much deeper than any other Cycle 1 program (to our knowledge, program 2079 being deepest with 17 hrs in F200W). We split the time roughly evenly between these two medium filters, as we do not know at what redshift a candidate might appear.
For cases around , F182M may only be a partial dropout, and we must demonstrate a lack of flux shortward of 1.7 m. The existing F150W data would do this, but not as deeply as this 2 m data warrants. We therefore add depth in F162M. Because the false positive interlopers are mid- galaxies with steeply falling SEDs, it is more efficient to test the dropout in a narrower band at the red edge of F150W: this includes all of the hypothesized signal and less background noise. This same argument is why it is better to observe F182M & F210M, rather than F182M & F200W: once one has found no flux in F182M, it is more efficient to seek flux in the redder half of F200W with the medium-band filter.
For the matching time in the LW arm, the key filter is F250M, through which one is trying to determine the continuum just longward of a Ly drop or detect a yet-higher redshift dropout in F250M itself. But one must guard against mid-redshift emission-line contamination in F250M, and for this we observe F300M and F335M. F300M is just short enough that if [OIII] 5007 is in F250M, then H will miss F300M and be found in F335M. This pairing therefore gives a robust measurement of the continuum.
By placing these observations on the parallel field created by program 1210, these new data build on the 55 hours of imaging depth already invested here. Given the steep number counts of galaxies at the redshift frontier, it is optimal to maximize the detection rate by integrating deeply on a single NIRCam pointing. The region in and around the HUDF in GOODS-S has the deepest and most comprehensive imaging and spectroscopy on the sky. While galaxies could be found anywhere, we site this program in GOODS-S to utilize and build upon the amazing legacy value of this field, while utilizing a position angle that allows ultra-deep MSA spectroscopy on the HUDF at a time of low zodiacal background.
3.4 Coordinated Parallels with NIRSpec
| Exposure Times (ksec) | |||||||||
| Subsurvey | Program | # Fields | Subpointings | # Targets | Prism | G140M | G235M | G395M | G395H |
| GOODS-S Deep/HST | 1210 | 1 | 3 | 250 | 100 | 25 | 25 | 25 | 25 |
| GOODS-S Year 2 | 3215 | 1 | 5 | 228 | 168 | 42 | 168 | ||
Program 3215 follows the path of 1210, using the NIRSpec MSA as the prime instrument and NIRCam as the coordinated parallel. This results in exceptionally deep spectroscopy on the HUDF and JADES Deep mosaic, about 104 hrs of total exposure, following on the 55 hrs of program 1210 spectroscopy released in Bunker et al. (2023a).
Following the program goal of exploring the high-redshift universe, we optimize this time toward the study of high-redshift galaxies. Unlike program 1210, which was largely targeted from HST ACS optical and WFC3 infrared imaging (Bunker et al., 2023a) save for four early high-redshift candidates from JADES (Robertson et al., 2023; Curtis-Lake et al., 2023), program 3215 will largely utilize NIRCam-selected targets. This allows a great expansion in the number of targets, as the JADES data is comfortably deeper than even the deepest WFC3 HUDF data and much deeper than the flanking CANDELS imaging.
Bringing ultra-deep spectroscopy to these faint high-redshift galaxies addresses many windows to understanding the physical processes associated with the early galaxy formation. We split the spectroscopic time across three dispersing modes—up to 168 ks on source with G395M (far deeper than program 1210), 168 ks with the Prism, and 42 ks with G140M—to serve a variety of key science drivers in studies of these galaxies.
The choice of obtaining ultra-deep spectra in G395M band comes from the requirement of probing the primary rest-optical lines at , with H down to . For galaxies at , the strong [OIII] 5007,4957Å doublet and H line are available to NIRSpec. H is available at . As has been seen in JADES and other surveys, the medium-resolution grating (G395M/F290LP) is excellent for studying these lines, producing precise redshifts, splitting closer pairs, and revealing kinematic signatures and broad AGN Balmer lines.
The deep Prism spectra are needed to determine the redshift for faint galaxies at through the detection of the Ly drop. The prism also can reveal rest-UV lines if they are strong enough, and generally provides a powerful route to measure the line flux of isolated lines over the full NIRSpec wavelength range.
Finally, we seek to characterize the role of neutral hydrogen in and around the galaxies in the reionization epoch (e.g. Jones et al., 2023a; Saxena et al., 2023; Witstok et al., 2023) by measuring the Ly emission line. The PRISM resolution is particularly coarse at the blue end of the spectrum, such that the grating substantially outperforms it for the detection of narrow lines. Further, because of the low resolution, the prism does not distinguish well between mild Ly emission and the Ly damping wing. The medium-deep G140M/F070LP spectra are therefore required and adequate to measure the Ly narrow line better than the prism and protect the damping wing analyses.
The ultradeep spectra delivered by this program will be uniquely suited to pursue several key science cases, as discussed in the following.
The G395M data will map the early mass-metallicity relation (MZR) at , reaching the unexplored low-mass regime at . Indeed, although major progress has been made with JWST in measuring the MZR and the Fundamental Metallicity Relation at z5 (e.g. Curti et al., 2023b, a; Nakajima et al., 2023), the low mass regime remains still a largely uncharted territory, except for rare, highly lensed cases (Vanzella et al., 2023) or cases where masses are not yet robustly determined (Maseda et al., 2023). With the ultra-deep spectra, we expect to accurately measure the metallicity via the detection of auroral lines, such as [OIII]4363 (Laseter et al., 2023; Sanders et al., 2023a) for tens of galaxies, and via strong line diagnostics (re-calibrated with the auroral-detected spectra) for fainter galaxies, possibly revealing signatures of Population III (Nakajima & Maiolino, 2022; Wang et al., 2022; Maiolino et al., 2023b). The spectra are also expected to provide good detections of the CIII]1907,1909 and OIII]1661,1666 transitions, which will be used to determine the C/O abundance ratio, which (due to the different enrichment timescales of carbon and oxygen) is an important clock for the galaxy star formation history (e.g. Maiolino & Mannucci, 2019; Arellano-Córdova et al., 2022; Jones et al., 2023b).
Recent results, based on a few tens of galaxies, have shown that the gas excitation increases steeply at . This trend is likely a consequence of higher ionization parameter, lower metallicity, and a harder radiation field, possibly resulting from stellar -enhancement or contribution from X-ray binaries, or even an excess of cosmic rays (Trump et al., 2023; Cameron et al., 2023; Fujimoto et al., 2023b; Katz et al., 2023; Sanders et al., 2023b; Trussler et al., 2022). The 3215 deep spectra will explore these trends towards lower mass galaxies, constraining the physics of the ISM with unprecedented accuracy at these redshifts.
In addition to providing spectroscopic confirmation of galaxies at the redshift frontier, deep prism spectra of distant galaxies have already illustrated the capability of providing tight constraints on stellar population parameters, finding very young ages (a few tens of Myr), very low metallicities (down to ), high yields of ionizing photons, low dust attenuation (Curtis-Lake et al., 2023; Atek et al., 2023). Yet, some of these constraints are still highly uncertain and often limited to relatively massive galaxies. These ultra-deep spectra will provide much tighter constraints on these parameters and push to much lower masses. For about 50 galaxies at z5 with S/N10 per resolution element on the continuum, it will even be possible to derive star formations histories (SFHs Looser et al., 2023a, b), hence testing the highly stochastic behavior of primeval galaxies expected by some models (Ceverino et al., 2018).
JWST spectroscopy has opened the exploration of the realm of intermediate mass black holes at high redshift, by detecting tens of AGN hosting black holes in the mass range between and , well below the BH mass range probed by quasar surveys (Furtak et al., 2023; Greene et al., 2023; Kocevski et al., 2023; Kokorev et al., 2023; Maiolino et al., 2023a; Matthee et al., 2023b; Übler et al., 2023). These studies have detected BHs in AGN that are still relatively luminous ( erg/s), as they were limited by the capability of detecting the nebular lines that can distinguish AGN from star forming galaxies. The ultra-deep spectra in program 3215 will allow the detection of even weaker active nuclei in early galaxies. Such AGNs can be recognized by their high ionization emission lines (NeV, NeIV, HeII), as well as other strong line diagnostics (Brinchmann, 2023; Nakajima & Maiolino, 2022; Tozzi et al., 2023). Further, in unobscured cases, we expect to detect the broad component of the H and H lines in grating spectra, separating this from gas outflows by comparing to the strong [OIII] lines. Comparing the inferred black hole masses to the host galaxy stellar mass will shed light on the formation channel of the first generation of black holes, differentiating between scenarios such as direct collapse, merging nuclear clusters, and Population III remnants (Habouzit et al., 2022; Volonteri et al., 2022; Pacucci et al., 2023; Schneider et al., 2023; Trinca et al., 2023).
Target selection for the program 3215 spectroscopy will be described fully in a later paper. In brief, we have prioritized high-redshift candidates from JADES, progressively working to lower redshift. We include rare special targets at high priority, incorporating AGN indicators, quiescent galaxies at high redshift, sub-mm detections, and strong line emitters extracted from slitless spectroscopy from the FRESCO survey (Oesch et al., 2023). As reported in Table 2, we could allocate shutters to 228 unique targets. These are slightly fewer targets than allocated in program 1210, because in program 3215 we have given priority in re-observing the same targets in the five pointings, so to maximize the total exposure time for each target.
Figure 4 shows the distribution of redshift and F444W magnitude for the targets that have been allocated shutters in the NIRSpec observation, illustrating that a significant number of them are very faint (). The color coding gives the total exposure time on each target (for the prism configuration), showing that many targets are observed for the full integration time of 168 ksec by being re-allocated shutters in the five pointings. We note that the three highest redshift galaxies from Curtis-Lake et al. (2023) are targeted for re-observation.
3.5 Additional Legacy Science
In addition to the primary science of probing the uncharted regime and revealing galaxy physics with ultra-deep spectroscopy in the parallel observations, the JOF data will provide a rich legacy data set of medium-band imaging and ultra-deep spectroscopy. We highlight here some examples of these opportunities.
The deep 1.6–4.1 m medium-band imaging will allow robust identification of quiescent galaxies via their Balmer breaks down to stellar mass. These data will constrain the evolution of the quiescent mass function and measure morphological gradients in Balmer break strengths (Ji et al., 2023), providing crucial constraints on the quenching mechanisms (e.g., Carnall et al., 2023; Nanayakkara et al., 2022; Santini et al., 2022; Marchesini et al., 2023; Looser et al., 2023a, b). The exquisite depth will also probe dwarf neighbors of larger mid-redshift galaxies, permitting analysis of their stellar populations and star-formation rates (Suess et al., 2023).
Turning to star-forming galaxies, by comparing the short-time-scale star-formation rates inferred from the measurement of H emission in the medium-bands to the longer time-scale values inferred from the rest-UV continuum, these data can test whether burstiness increases out to (e.g., Emami et al., 2019; Faisst et al., 2019; Reddy et al., 2023; Looser et al., 2023b). Bursts are hypothesized to be a critical driver of the strong emission lines observed at high- (e.g., Stark et al., 2013; Faucher-Giguère, 2018). With this very deep imaging, one can also map the line emission and hence star-formation rates across the face of these galaxies.
The data will also be an opportunity for the infrared time-domain, as the return to the deep Cycle 1 field with very deep year 2 imaging will reveal variability at micron on a timescale of year with medium-band resolution. This will probe supernovae (SNe) at (and potentially beyond), for which the SN emission shifts to micron with a peak brightness down to AB mag, as well as detect some AGN by their photometric variability.
3.6 Implementation
Program 3215 is implemented in the same manner as 1210, described in Eisenstein et al. (2023). NIRSpec must be the prime instrument with NIRCam in coordinated parallel, but by using multiple MSA configurations at slightly different pointings, we can achieve a 15-point dither pattern for NIRCam.
We split the program into 5 visits, each with a pair of co-pointed MSA designs, one for the prism and one for the grating. Each visit contains 9 disperser/filter pairs, and each of these is observed for a 3-shutter nod. Each exposure contains 2 integrations for NIRSpec (19 groups of NRSIRS2 readout, Rauscher et al. 2017, 1400 s each) as well as 2 integrations for NIRCam (7 groups of DEEP8 readout, 1375 s each).
In each of the 5 visits, we perform 4 nods with the G395M, 1 nod with G140M, and finally 4 nods with the PRISM. These are paired with 6 NIRCam exposures of F162M & F335M, 12 exposures of F182M & F250M, and 9 exposures of F210M & F300M. The total NIRCam exposure times are summarized in Table 1.
The exact pointings of the 5 MSA pairs are chosen to place a small number of candidates onto all of the designs, while providing pixel diversity for both NIRCam and NIRSpec. We keep the total spacing small, under 2, as this increases the number of targets that can be observed on legal shutters in most of the designs. We also use the same position angle as 1210 (V3 PA of 321) and minimize the displacement of the two programs, to ensure high overlap of 14 filters.
As of this writing, program 3215 is scheduled to be observed in late October 2023, essentially on the one-year anniversary of the Year 1 observations of programs 1210 and 1180 in this field. Because program 3215 has no proprietary period, these data sets will go public at nearly the same time.
4 Cycle 3: NIRCam Grism Spectroscopy with Program 4540
4.1 Science Motivation
While the investment of JWST imaging time in the JOF is tremendous, this field does not yet have the spectroscopic resources of the HUDF. The JOF lies in the flanking depth region of GOODS-S and CANDELS, rather than in the HUDF itself. GOODS-S has had substantial spectroscopic attention from ground-based observatories and the 3D-HST survey, but it does not have the ultra-deep data that the HUDF, particularly from the deep VLT/MUSE program. The JOF also has only limited overlap with the JWST FRESCO program. In December 2023, this area will be targeted with 4-5 NIRSpec MSA configurations from program 1286 and 1287, which will provide coverage of some high-priority objects.
The NIRCam grism has been shown to produce a compelling capability to measure strong [OIII] and H emission lines from galaxies at (Sun et al., 2023a; Kashino et al., 2023; Matthee et al., 2023a; Oesch et al., 2023; Wang et al., 2023c). Such observations would be an excellent pairing to the imaging, providing spectroscopic redshifts, line diagnostics, and dynamical measurements to this deep window on the universe. The NIRCam instrument team therefore has decided to devote 32 hours of Cycle 3 guaranteed time toward such a program (program ID 4540, PI: Eisenstein).
There are a number of key motivations from the science opportunity from these data. Some of these are generic to NIRCam grism data, but many are specific to the opportunities of this special field.
The core goal is to make make more complete use of the opportunity at that the [OIII] and H lines are often extremely strong. Observing only one filter leaves a redshift gap, but observing in both the F356W and F444W filters provides complete coverage at with these two strong lines. Further, at –6.6, the data provide access to both lines, which is useful for studying the metallicity of the galaxy populations and which nails the line identification.
At –6.6, these data will include both H and H, the ratio of which allows us to infer the amount of dust reddening, indeed free from the complicated path-loss corrections that result from the NIRSpec MSA.
At –5.2, we will measure H in the grism but will also measure [OIII] and H in the F250M, F300M, and F277W imaging. This will allow us to extend the measurement of the ratio of [OIII] to Balmer lines to a larger population of galaxies.
These same strong lines are usually detected in the medium-band imaging, where comparison between the filters allows one to measure the morphology in the line, as distinct from the rest-optical continuum. When combined with the grism data at a spectral resolution , one can break the degeneracy between the velocity field and the morphology, providing spatially resolved velocity measurements in some galaxies (Nelson et al., 2023).
Extending this opportunity, we note that a portion of these program 4540 observations overlap cycle 1 F444W grism observations from FRESCO (Oesch et al., 2023) but with a dispersion direction that differs by about 50. This will improve the velocity field measurements, as slitless spectroscopy always has a degeneracy when there is a velocity gradient across the object in the dispersion direction. It also provides another way to confirm the association of grism lines to the galaxies in the undispersed images.
The high spectral resolution of the NIRCam grism yields important dynamical information from the line widths of the emission lines. The widest lines can indicate galactic outflows or AGN broad lines, giving critical clues to galactic feedback and the role of black holes in galaxy growth (Matthee et al., 2023b; Maiolino et al., 2023a; Greene et al., 2023; Furtak et al., 2023; Harikane et al., 2023b; Kokorev et al., 2023).
The spectroscopy also creates important opportunities for the modeling of the detailed infrared SEDs. For , these spectroscopically detected galaxies will have their Balmer breaks in the progression of medium bands from program 3215; this will allow careful characterization of star formation histories. The detailed SED shapes around the Balmer and 4000Å breaks are key for detecting stars older than about 50 million years, but the effects are subtle and can be confused by degeneracies between the rest-frame colors and the photometric redshifts. Limits on the strong Balmer lines also allow one to clean the effects of the weaker Balmer lines around 4000Å from the continuum photometry. Finally, the strong lines themselves can be included in the galaxy spectral synthesis analyses, as a constraint on the most recent star formation.
Some of these galaxies will have very deep MIRI F770W data from program 1180, probing rest-frame 1–2 at these redshifts. This will provide more robust constraints on stellar populations and in particular on the stellar mass of these emission-line galaxies. We stress the opportunity to compare these stellar masses to the dynamical masses from resolved velocity profiles. Although the MIRI data will only detect a subset of the galaxies, that subset can be used to refine the analysis in the NIRCam bands to improve mass determinations overall.
Moving beyond the modeling of single galaxies, an exciting aspect of the NIRCam grism is the ability to detect a rich pattern of galaxy overdensities at these high redshifts. The FRESCO, EIGER, and ASPIRE data sets have shown beautiful results in this regard (Kashino et al., 2023; Witstok et al., 2023; Helton et al., 2023; Wang et al., 2023c; Herard-Demanche et al., 2023; Sun et al., 2023b). The patterns are much easier to find in three-dimensional mapping, and the ability of the grism to observe nearly every galaxy in an unbiased way has been crucial for getting to dense enough coverage to show these structures. In doing so, we can place the galaxies in their environmental context, connecting all of the imaging information about morphology, SEDs, and emission-line properties to environment. We judged this opportunity to measure this large-scale structure context so compelling that we opted to spread this grism data to cover the inter-module gap, even though programs 1210 and 3215 do not cover it, so that we can measure the environment of galaxies that fall near the edge of the imaging modules.
Indeed, Helton et al. (2023) presents the discovery of a substantial arcminute-scale galaxy overdensity with 43 emission-line galaxy members extracted from the south-west corner of the FRESCO data. Some of the galaxies fall exactly in the program 4540 footprint, which will extend the mapping beyond the FRESCO boundary. Our deeper data will better measure the dynamics of this forming structure through the [OIII] and H detections of these galaxies.
Finally, while the above focuses on the strong [OIII] and H lines, these grism data will net spectroscopic detections at lower redshift of the weaker rest-infrared lines, confirming redshifts to strong star-forming dwarfs and measuring lines whose equivalent widths are likely too small to claim from the imaging. We particularly note the ability to detect Paschen to (and Paschen to ), the comparison of which to the rest-UV continuum measured in the HST imaging in this field and H available in the F150W and F162M (and out to F250M, for P) filters can more directly reveal dust-obscured star formation. He I 10830Å and [S III] 9532Å lines can also be detected in vigorous cases, with line identification coming from the medium-band detections of the stronger lines at bluer wavelengths.
4.2 Short-wave Data
We use the opportunity of program 4540 to add deep F070W imaging to the JOF, the final NIRCam broad-band filter and 15th NIRCam filter overall. Given the mosaic pattern described below, this will reach exposure times of 16–68 ks, depending on overlaps, mildly shallower than but comparable to the 55 and 73 ks in F090W and F115W, respectively. This will extend the JWST coverage of the SEDs and produce optical imaging much deeper than the ACS imaging from GOODS and CANDELS.
These F070W data provide key enhancements for the core science of the program, as it will provide direct sensitivity to a fainter population of Ly break galaxies at –6. This will provide a UV-selected population to be compared to the emission-line selected population in [OIII] and H and to the characterization of Balmer/4000Å breaks from the medium-band data.
4.3 Implementation
Driven by the desire to cover most of the 3.1 to 5 m range over the full program 1210 footprint as well as to cover the intermodule gap and a small boundary for environment measurements, we design program 4540 for a tight 22 mosaic pattern as shown in Figure 5. The same data will give partial wavelength coverage in a mildly broader area; this bounding area is largely covered by JADES medium-depth NIRCam imaging.
As redshifts in this field are so valuable, we aim to achieve a depth that is notably fainter than the FRESCO observations. In the overlap regions of the 4 tight pointings, we will reach a depth of over 9 hrs of integration time in each filter, compared to 2 hrs in FRESCO F444W. The F356W filter data is yet more sensitive because of lower zodiacal background. We expect to measure several hundred redshifts at , in addition to the lower redshift galaxies detected in lower luminosity redder lines.
The grism data are undeflected at 3.95 , so the regions of complete F356W spectra and complete F444W spectra fall in different portions of the NIRCam field of view. With the column-direction grism (GRISMC), we use 2 rows of the mosaic stepped so that we can get complete coverage over the full JOF NIRCam imaging footprint. We then offset between F356W and F444W mosaics by so that we get complete 3.1 to 5 coverage over a consistent footprint. GRISMC is preferred to GRISMR for this purpose because the dispersion direction is the same in the two modules.
Following the pattern well-used in our earlier GTO program, we are using the DEEP8 readout mode to minimize data volume. We will use long integrations (6 or 7 groups, 1159–1375 sec) to minimize overheads and reduce the impact of detector noise. We use 6–8 dither positions in each of the mosaic positions to produce good pixel redundancy and achieve the desired exposure time The INTRAMODULEBOX pattern is tightest and most optimal for our purpose. Because of allocation constraints, we have slightly different exposure times per pointing, and we use longer exposures on the two F356W grism images with more coverage of the JOF.
Although slitless spectroscopy often uses multiple dispersion directions to robustly connect lines to the zero-order location and hence get unique wavelengths, we do not do this here. With the sharp NIRCam imaging and accurate grism spectral tracing models, the associations of lines to galaxies detected in the imaging is often unconfused. More importantly, this field has exceptionally deep multi-band imaging, and these high-equivalent emission lines (essentially all of the cases) show up easily in the medium-bands from programs 1180, 1210, and 3215. This sharply reduces concerns about the ability to robustly match up the line detections to the undispersed images. Further, the NIRCam imaging yields very high quality photometric redshifts, which reduces the pressure to have a second line detection to produce an identification of the line and hence a redshift.
Direct images are required by the slitless spectroscopy observing template. We use short exposures (5–6 groups of SHALLOW4, 257–311 sec) because this field already has far deeper data. We use the same LW filters, but change in SW to F150W and F200W to increase the signal-to-noise in the short exposures and to assure some infrared coverage in the outer portions of field where we may fall off of previous NIRCam imaging.
In detail, we found that layout of this program in a rigid mosaic was impossible given a hole in guide star coverage. Because of the two out-of-field direct images per pointing and the offset between the pointings of the two filters, there are a total of 24 separate visits. To achieve schedulability at a consistent position angle, we had to decouple the 8 grism positions and make small (10-20) level adjustments to find guide stars. Hence, the mosaic is an irregular grid.
4.4 Coordinated Parallels with MIRI
Program 4540 will also produce coordinated parallels for MIRI imaging in the F770W and F1500W filters. For each filter, these will reach a depth of about 2 hrs in each of 4 overlapping pointings, with a peak exposure time around 9 hrs. We choose to use F770W in the outermost 4 pointings, yielding a contiguous field of about by , while we use F1500W in the inner 4 pointings, yielding a field of by that is contained within the F770W mosaic. As with JADES, we use SLOWR1 readout mode to limit the data volume, and we use multiple integrations per NIRCam integration to mitigate saturation.
These F770W data will extend the study of galaxies and particularly their stellar mass utilizing rest-frame 1–2 measurements, while the the F1500W data will provide insight into the role of AGN in –5 galaxies, as the hot dust in the nuclei can outshine the Rayleigh-Jeans tail of the stellar population. Finding such sources may be the most powerful means to identify heavily obscured AGN at .
However, the location of this parallel field does fall into an area currently lacking in NIRCam and HST imaging. Of course, the extended CDF-S does have deep ground-based imaging (such as the deep 5-year stack from the Dark Energy Survey SNe program) as well as imaging from Spitzer, which can support near-term use. We expect that this field is likely to be imaged with NIRCam in the future, as its location is easy to reach as a parallel when NIRSpec is pointed at the exceptional JOF NIRCam imaging, indeed at the position angle such that the NIRSpec quadrants align to the NIRCam geometry.
5 JADES Second Imaging Data Release
In this paper, we present the second imaging data release of JADES, including the GOODS-S imaging observed in September and October 2022. The first release (Rieke et al., 2023) included the NIRCam imaging of the Deep Prime program on the HUDF. This second release includes the deep NIRCam parallel from program 1210 and the medium-depth mosaic that flanks these two imaging fields. Caveats regarding data quality and some missing data due to short circuits in the NIRSpec MSA and one skipped visit are described in Eisenstein et al. (2023).
Figure 6 shows a three-color image of the field. The total area of the JADES imaging (excluding the FRESCO-only region) is 67.7 arcmin, and the catalogs contain 94,000 distinct objects. A summary of the image reduction and catalog processing follows.
To create calibrated imaging and object photometry catalogs, the NIRCam data are reduced in the same manner as described in Rieke et al. (2023, hereafter R23) with a few modifications. For this release, version 1.11.4 of the STScI jwst pipeline is used but with a custom astrometric system as discussed in R23. The JWST calibration reference data file used during processing was jwst_1130.pmap, which adjusted the photometric zeropoints slightly relative to the first JADES data release. We have checked that the flux ratios of objects shared between the current data release and Rieke et al. (2023) are broadly consistent with the changes in photometric zeropoint.
To increase our pixel redundancy, the JADES V2.0 mosaics are constructed by simultaneously stacking all available exposures at each location rather than the visit-based combination and stacking used in R23. To achieve this without taxing computer memory, we split each mosaic into smaller sub regions that can be processed separately. For F090W, F115W, and F150W, we remove a small number of hot pixel residuals from the pipeline by replacing the affected pixels with the corresponding pixels from median-filtered versions of the mosaics. After visual inspection of individual exposures, we masked regions of high persistence or with stray light effects with high spatial frequency, especially noticeable in the 1180 Medium region. This masking reduces the depth and exposure time in only a small fraction of the total JADES area.
The exposure times achieved in this mosaic are shown in Figure 7. These show the deep parallel field of program 1210, along with the year 1 data in the Deep and Medium mosaics of program 1180. Year 2 observations will double the depth of the Deep mosaic, add flanking Medium depth around the Deep region, and extend the footprint with parallel NIRCam imaging from program 1286 and 1287 (Eisenstein et al., 2023). Figure 8 shows maps of the corresponding 5 sensitivity for F115W, F200W, and F356W in diameter apertures, with aperture corrections applied. These maps are computed by placing randomized apertures throughout the fully reduced mosaic images, after masking sources and surrounding pixels with significant flux. As reported by Rieke et al. (2023), the 1180 Deep region reaches F200W depth of mag AB.
To enhance the multi-band space-based photometry available for sources, we also perform photometry on the Hubble Legacy Field HST/ACS F435W, F606W, F775W, and F814W mosaics and HST/WFC3 F105W, F125W, F140W, and F160W mosaics (Illingworth et al., 2016) after rectifying them to our astrometric frame. We reprocess and release JWST/NIRCam F182M and F210M mosaics from data acquired by the JWST Extragalactic Medium-band Survey (JEMS; Williams et al., 2023) and First Reionization Spectroscopically Complete Observations (FRESCO; Oesch et al., 2023) programs. We additionally reprocess and release JWST/NIRCam F430M, F460M, and F480M images from data acquired by JEMS and combine reprocessed JWST/NIRCam F444W exposures originally acquired by FRESCO with our F444W stack. We include photometry from all JWST filters in our released catalogs, where available.
The construction of the photometric catalogs largely follows that described by Rieke et al. (2023). The LW channel images are stacked to create an inverse-variance-weighted signal-to-noise ratio image for performing detection. Catalogs are constructed from the resulting segmentations, with automated and manual masking of stars and diffraction spikes applied. SExtractor (Bertin & Arnouts, 1996) windowed positions are used as object centroids. Aperture photometry is measured using a range of circular apertures, two Kron apertures, and segmentations as discussed in Rieke et al. (2023). Sky background uncertainties are computed by placing random apertures throughout the images and directly from the jwst pipeline ERR image layers when available. Aperture corrections are computed from model point-spread-functions constructed following Ji et al. (2023). Flags are set to indicate bad pixels that may impact object segments, whether sources may be influenced by brighter neighbors, and whether stars identified in the image may impact them or if the objects themselves are consistent with stars.
In addition to the photometry, we have computed photometric redshifts following the methods described in Rieke et al. (2023) and Hainline et al. (2023). In particular, these use the EAZY code (Brammer et al., 2008) with templates as described in Hainline et al. (2023) and incorporating percent-level adjustments to the photometric zeropoints based on the global median offset of each filter in a first-round of fitting. For convenience, the photometric redshift catalog is included as FITS extension of the photometry catalog.
The images and catalogs are available on the Mikulski Archive for Space Telescopes (MAST) at https://archive.stsci.edu/hlsp/jades, with DOI 10.17909/z2gw-mk31. We will incorporate the 3215 imaging using this same pipeline in an incremental release.
6 Discussion
We have described the executed and approved plans for JWST observations of the JADES Origins Field, spanning the first three mission cycles. In total, the time to be invested in this single NIRCam footprint of 9 arcmin is already very large, roughly 360 open-shutter hours of exposure time across in 15 NIRCam imaging filters, 20 hours of NIRCam slitless spectroscopy, and over 170 hours of MIRI imaging. This is the most intensively observed JWST field yet, designed to probe the redshift frontier beyond .
Figure 9 shows an small portion of the JOF imaging data from program 1210 to highlight the opportunities of the deep NIRCam medium-band data, here F335M and F410M. Strong emission lines are revealed easily in these images, with beautiful morphological detail. This reach will be greatly extended in program 3215 with 5 new medium bands spanning 1.6 to 3.0 .
The location of the JOF in close proximity to the HUDF and the Deep Prime JADES NIRCam imaging mosaic adds considerable value to the GOODS-S region. Wider field telescopes can capitalize on both regions, and the parallel opportunities will continue to appear with JWST. By construction, further NIRSpec observations of the HUDF can place further NIRCam exposure time in the JOF region; the position angle is at a period of low zodiacal background and is allowable with the micrometeorite constraint. Indeed, program 3215 is using this opportunity to conduct exceptionally deep spectroscopy (104 hrs of total exposure) in the HUDF, focusing on the physics of high-redshift galaxies. Further, observations with NIRISS or MIRI on the JOF are easily paired with NIRCam observations on and around the HUDF, within the JADES Deep NIRCam mosaic.
Accompanying this paper, we have released the reduced images and catalogs of the Year 1 JADES data in and around the JOF. These data are already among the very deepest observed with JWST NIRCam, with over 110 hrs of total exposure spread across 9 filters. These data are available at https://archive.stsci.edu/hlsp/jades and a feature-rich FITSmap viewer is provided at https://jades.idies.jhu.edu, with a version tailored for the public at https://jades.idies.jhu.edu/public. Observations of the Cycle 2 program 3215 are underway at the time of this writing (Oct. 2023) and will be immediately public.
The power of JWST to reveal the early history of galaxies with infrared imaging and spectroscopy is both transformative and inspiring. Many compelling observing programs have been performed or are underway, each with their own nuances and opportunities. We hope that the extraordinary depth and filter complement of the JADES Origins Field will provide a superb new view of galaxies in their youth.

### v2

@@
 Abstract
-We summarize the properties and initial data release of the JADES Origins Field (JOF), which will soon be the deepest imaging field yet observed with the James Webb Space Telescope (JWST).
+We summarize the properties and initial data release of the JADES Origins Field (JOF), the longest single pointing yet imaged with the James Webb Space Telescope (JWST).
 This field falls within the GOODS-S region about 8’ south-west of the Hubble Ultra Deep Field (HUDF), where it was formed initially in Cycle 1 as a parallel field of HUDF spectroscopic observations within the JWST Advanced Deep Extragalactic Survey (JADES).
-This imaging will be greatly extended in Cycle 2 program 3215, which will observe the JOF for 5 days in six medium-band filters, seeking robust candidates for galaxies.
-This program will also include ultra-deep parallel NIRSpec spectroscopy (up to 104 hours on-source, summing over the dispersion modes) on the HUDF.
-Cycle 3 observations from program 4540 will add 20 hours of NIRCam slitless spectroscopy to the JOF.
-With these three campaigns, the JOF will be observed for 380 open-shutter hours with NIRCam using 15 imaging filters and 2 grism bandpasses.
+This imaging was greatly extended in Cycle 2 program 3215, which observed the JOF for 5 days in six medium-band filters, seeking robust candidates for galaxies.
+This program also includes ultra-deep parallel NIRSpec spectroscopy (up to 91 hours on-source, summing over the dispersion modes) on the HUDF.
+Cycle 3 observations from program 4540 added 20 hours of NIRCam slitless spectroscopy and F070W imaging to the JOF.
+With these three campaigns, the JOF was observed for 380 open-shutter hours with NIRCam using 15 imaging filters and 2 grism bandpasses.
 Further, parts of the JOF have deep 43 hr MIRI observations in F770W.
-Taken together, the JOF will soon be one of the most compelling deep fields available with JWST and a powerful window into the early Universe.
+Taken together, the JOF is one of the most compelling deep fields available with JWST and a powerful window into the early Universe.
 This paper presents the second data release from JADES, featuring the imaging and catalogs from the year 1 JOF observations.
+ma
 1 Introduction
@@
 Precisely timing when the first galaxies formed after the Big Bang is a key goal of observational cosmology and one of the last great unknowns in our history of the Universe: When and how did the first galaxies form?
-The primordial star formation and the formation of galaxies in the first 300 Myr (redshift ) after the Big Bang is a sensitive probe of the physics governing the conversion of gas into stars at early times and different dark matter models (e.g., Behroozi & Silk, 2015; Dayal & Ferrara, 2018; Boylan-Kolchin, 2023; Kannan et al., 2023).
-Additionally, these first, unexplored galaxies might harbor surprises such as metal-free stellar populations (Population III) or rapidly accreting young black hole seeds, thereby shedding light onto possible pathways to the formation of the billion-solar-mass black holes observed at (e.g., Bromm & Yoshida, 2011; Xu et al., 2013; Smith & Bromm, 2019; Vikaeus et al., 2022; Volonteri et al., 2021; Trinca et al., 2023; Schneider et al., 2023).
+The primordial star formation and the formation of galaxies in the first 300 Myr (redshift ) after the Big Bang is a sensitive probe of the physics governing the conversion of gas into stars at early times and different dark matter models (e.g., P.
+S.
+Behroozi & J.
+Silk, 2015; P.
+Dayal & A.
+Ferrara, 2018; M.
+Boylan-Kolchin, 2023; R.
+Kannan et al., 2023).
+Additionally, these first, unexplored galaxies might harbor surprises such as metal-free stellar populations (Population III) or rapidly accreting young black hole seeds, thereby shedding light onto possible pathways to the formation of the billion-solar-mass black holes observed at (e.g., V.
+Bromm & N.
+Yoshida, 2011; H.
+Xu et al., 2013; A.
+Smith & V.
+Bromm, 2019; A.
+Vikaeus et al., 2022; M.
+Volonteri et al., 2021; A.
+Trinca et al., 2023; R.
+Schneider et al., 2023).
 Deep fields, both in imaging and associated spectroscopy, are a key opportunity for this work, allowing the astronomy community to coordinate its multi-wavelength resources to best advantage.
 Cycle 1 observations with JWST succeeded quickly in pushing the redshift frontier beyond what had been seen with HST.
-Relying on the complete scattering of photons shortward of Ly 1216Å by the intergalactic medium (IGM), many F115W dropouts at and even some F150W dropouts at were detected as photometric candidates in numerous fields (e.g., Bouwens et al., 2022; Castellano et al., 2022; Finkelstein et al., 2022; Harikane et al., 2023a; Naidu et al., 2022b; Adams et al., 2023; Austin et al., 2023; Atek et al., 2023; Donnan et al., 2023; Franco et al., 2023; Hainline et al., 2023; Wang et al., 2023a; Whitler et al., 2023; Yan et al., 2023).
-As of this writing, numerous galaxies at , including two at , have been spectroscopically confirmed with long NIRSpec exposures (Curtis-Lake et al., 2023; Heintz et al., 2023; Morishita & Stiavelli, 2023; Roberts-Borsani et al., 2023; Tang et al., 2023; Arrabal Haro et al., 2023b; Bunker et al., 2023b; Hsiao et al., 2023; Sanders et al., 2023b; Wang et al., 2023b; Fujimoto et al., 2023b), along with a larger number of galaxies out to where the much brighter [OIII] and H lines are available in the NIRSpec bands.
-These early results have demonstrated not only JWST’s unrivalled capabilities in successfully exploring the early Universe, but also its potential for pushing the redshift frontier even further.
-Indeed, it is evident that JWST—with its excellent design and on-orbit performance (Rigby et al., 2022)—is fully capable of probing back to even earlier cosmic times.
+Relying on the complete scattering of photons shortward of Ly 1216Å by the intergalactic medium (IGM), many F115W dropouts at and even some F150W dropouts at were detected as photometric candidates in numerous fields (e.g., M.
+Castellano et al., 2022; S.
+L.
+Finkelstein et al., 2022, 2024; R.
+P.
+Naidu et al., 2022a; N.
+J.
+Adams et al., 2023; D.
+Austin et al., 2023; R.
+Bouwens et al., 2023; C.
+T.
+Donnan et al., 2023; Y.
+Harikane et al., 2023a; P.
+G.
+Pérez-González et al., 2023a; B.
+E.
+Robertson et al., 2023; L.
+Whitler et al., 2023; H.
+Yan et al., 2023; H.
+Atek et al., 2024; M.
+Franco et al., 2024; K.
+N.
+Hainline et al., 2024a).
+In the first year, numerous galaxies at , including two at , were spectroscopically confirmed with long NIRSpec exposures (E.
+Curtis-Lake et al., 2023; K.
+E.
+Heintz et al., 2023; T.
+Morishita & M.
+Stiavelli, 2023; G.
+Roberts-Borsani et al., 2023; P.
+Arrabal Haro et al., 2023a; A.
+J.
+Bunker et al., 2023; S.
+Fujimoto et al., 2023a; R.
+L.
+Sanders et al., 2023; M.
+Tang et al., 2023; B.
+Wang et al., 2023; C.
+M.
+Casey et al., 2024; T.
+Y.-Y.
+Hsiao et al., 2024; D.
+J.
+McLeod et al., 2024), along with a larger number of galaxies out to where the much brighter [OIII] and H lines are available in the NIRSpec bands.
+These early results demonstrated not only JWST’s unrivalled capabilities in successfully exploring the early Universe, but also its potential for pushing the redshift frontier even further.
+Indeed, it is evident that JWST—with its excellent design and on-orbit performance (J.
+Rigby et al., 2023)—is fully capable of probing back to even earlier cosmic times.
 Within this context, we note that the Cycle 1 programs were designed to go deep in F115W and F150W, not in F200W.
-The few hints of any F200W dropouts are uncertain, potentially mimicked by –5 galaxies with dust-reddened continua and well-placed strong rest-optical emission lines (Fujimoto et al., 2023a; Naidu et al., 2022a; Zavala et al., 2022; Arrabal Haro et al., 2023a).
-We describe here the construction of a new deep field, the JADES Origins Field (JOF), that aims to push our exploration of early galaxies further than ever before possible.
-This field was initiated by the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein et al., 2023) as JWST NIRCam coordinated parallel observations to a NIRSpec multi-object spectroscopy (MOS) pointing focused on the Hubble Ultra Deep Field (HUDF; Beckwith et al., 2006) within the GOODS-S field (Giavalisco et al., 2004).
-These Cycle 1 observations of JADES were described in Eisenstein et al.
-(2023) and Bunker et al.
-(2023a).
-New data in Cycle 2 and 3 will bring this field to nearly 10 mission days of NIRCam imaging and slitless spectroscopy in addition to over 7 mission days of deep MIRI imaging, making it the deepest JWST imaging yet achieved.
+The few hints of any F200W dropouts are uncertain, potentially mimicked by –5 galaxies with dust-reddened continua and well-placed strong rest-optical emission lines (S.
+Fujimoto et al., 2023b; R.
+P.
+Naidu et al., 2022b; J.
+A.
+Zavala et al., 2023; P.
+Arrabal Haro et al., 2023b; P.
+G.
+Pérez-González et al., 2023b).
+We describe here the construction of a new deep field, the JADES Origins Field (JOF), designed to push our exploration of early galaxies further than ever before possible.
+This field was initiated by the JWST Advanced Deep Extragalactic Survey (JADES; D.
+J.
+Eisenstein et al., 2023) as JWST NIRCam coordinated parallel observations to a NIRSpec multi-object spectroscopy (MOS) pointing focused on the Hubble Ultra Deep Field (HUDF; S.
+V.
+W.
+Beckwith et al., 2006) within the GOODS-S field (M.
+Giavalisco et al., 2004).
+These Cycle 1 observations of JADES were described in D.
+J.
+Eisenstein et al.
+(2023) and A.
+J.
+Bunker et al.
+(2024).
+New data in Cycle 2 and 3 brought this field to nearly 10 mission days of NIRCam imaging and slitless spectroscopy in addition to over 7 mission days of deep MIRI imaging, making it the deepest JWST imaging yet achieved.
 The geometry of this field is well suited for future JWST work, as the parallels can span between the JADES Deep mosaic around the HUDF and the JOF at observing windows that have low zodiacal background.
-As a start on this, NIRSpec MOS work in the JOF will be conducted as part of JADES in late 2023.
 In this paper, we detail the scientific design and implementation of the Cycles 2 and 3 observations.
@@
 As we argue in § 3 (and as mentioned above), broad-band NIRCam data alone is susceptible to selection of Balmer-break and emission-line interlopers from mid-redshift, and this failure mode appears to become more pernicious at , because of the rarity of these galaxies and because of the more limited number of applicable filters.
-Using medium-bands F162M, F182M, F210M, F250M, F300M, and F335M will exclude mid-redshift emission lines and allow for detection of more distinctive breaks.
-We expect that these data will provide robust observational constraints on the number density of galaxies and their rest-UV luminosities, colors, and sizes.
-Of course, these exceptionally deep images, up to 46 hours per filter and covering about 9 square arcminutes, will be of high value for the study of intermediate-redshift galaxies as well.
-To support this, the NIRCam instrument team will conduct NIRCam slitless spectroscopy in Cycle 3 in both the F356W and F444W filters, reaching up to 9 hr of on-source integration time.
+Using medium-bands F162M, F182M, F210M, F250M, F300M, and F335M can exclude mid-redshift emission lines and allow for detection of more distinctive breaks.
+These data provide robust observational constraints on the number density of galaxies and their rest-UV luminosities, colors, and sizes (B.
+Robertson et al., 2024)
+Of course, these exceptionally deep images, up to 46 hours per filter and covering about 9 square arcminutes, are of high value for the study of intermediate-redshift galaxies as well.
+JADES has conducted 5 NIRSpec MOS pointings on the JOF itself; these data will be presented in JADES Data Release 4 (Curtis-Lake et al., in prep; Scholtz et al., in prep.).
+An additional pointing using a novel ultra-high multiplex layout for the G235M and G395M grating was taken as part of program 3215 and will be presented in D’Eugenio et al.
+(in prep.).
+To augment this slit spectroscopy, the NIRCam instrument team conducted NIRCam slitless spectroscopy in Cycle 3 in the F322W2, F356W, and F444W filters, reaching up to 9 hr of on-source integration time.
 These will provide spectroscopic redshifts for hundreds of galaxies at with the strong H and/or [OIII] 5007Å lines.
 A further opportunity of the Cycle 1 and 2 data is the observation of a very deep NIRSpec MOS pointing on the UDF.
-The Cycle 1 data was already released in Bunker et al.
-(2023a); here we describe the science drivers and layout of the Cycle 2 pointing, which will target faint high-redshift galaxies with exposure times up to 104 hrs on-source.
+The Cycle 1 data was already released in A.
+J.
+Bunker et al.
+(2024); here we give an overview of the Cycle 2 pointing, which targeted faint high-redshift galaxies with exposure times up to 91 hrs on-source.
+Further description of the NIRSpec target selection and data properties is presented in JADES Data Release 3 (F.
+D’Eugenio et al., 2025a).
 In addition to an overview of the upcoming programs, this paper presents the initial release of NIRCam imaging data and catalogs on this field, already some of the deepest imaging obtained by JWST.
-The release includes medium-depth flanking fields that connect this footprint to the HUDF region released in the JADES first data release (Rieke et al., 2023).
+The release includes medium-depth flanking fields that connect this footprint to the HUDF region released in the JADES first data release (M.
+J.
+Rieke et al., 2023).
+Reductions of the deep imaging results will be in an upcoming data release.
 In § 2–4, we present the Cycle 1, 2, and 3 programs in the JOF. § 5 then presents the initial data release, and we summarize in § 6.
 2 Cycle 1: Beginning this field as a Coordinated Parallel
-The imaging in this field began as the NIRCam Coordinated Parallel of program 1210 (PI: Lützgendorf), which executed the JADES Deep/HST NIRSpec micro-shutter array (MSA) observations in and around the HUDF (Bunker et al., 2023a).
+The imaging in this field began as the NIRCam Coordinated Parallel of program 1210 (PI: Lützgendorf), which executed the JADES Deep/HST NIRSpec micro-shutter array (MSA) observations in and around the HUDF (A.
+J.
+Bunker et al., 2024).
 Fortuitously, the position angle of these observations put these NIRCam parallels on top of further data from program 1180 (PI: Eisenstein) that included medium-depth NIRCam imaging and deep MIRI F770W imaging, the latter resulting from coordinated parallels of deep NIRCam imaging around the HUDF.
-These observations are described in Eisenstein et al.
+These observations are described in D.
+J.
+Eisenstein et al.
 (2023).
@@
 The 1180 medium imaging overlaps and extends this geometry at flanking depth in all filters save F335M.
-The MIRI imaging will reach 43 hr depth in F770W in each of 4 adjoining fields.
-Further, an assigned observing window in Year 2 has allowed us to place one Coordinated Parallel NIRCam field from the JADES Medium/JWST tier (program 1286, PI: Lützgendorf) in a partially overlapping location about 1 south of the 1210 pointing.
+The MIRI imaging reach 43 hr depth in F770W in each of 4 adjoining fields.
+Further, an assigned observing window in Year 2 allowed us to place one Coordinated Parallel NIRCam field from the JADES Medium/JWST tier (program 1286, PI: Lützgendorf) in a partially overlapping location about 1′ south of the 1210 pointing.
 This is 11 hours open-shutter of additional data in all the previous nine filters plus F070W.
-It will add further depth to the JOF imaging, but also will make a deeper region (being the superposition of two medium surveys) that is more square and hence better suited for the NIRSpec MOS footprint.
+It adds further depth to the JOF imaging, but also makes a deeper region (being the superposition of two medium surveys) that is more square and hence better suited for the NIRSpec MOS footprint.
 The extra NIRCam imaging also overlaps more of the deep MIRI data.
-A simplified view of the GOODS-S region is provided in Figure 1, showing the location of the JOF relative to the HUDF, MUSE, and other JWST programs.
+A simplified view of the GOODS-S region is provided in Figure 1, showing the location of the JOF relative to the HUDF, CANDELS, and other JWST programs.
 3 Cycle 2: Deep Medium-Band Imaging with Program 3215
@@
 With the advent of JWST, we now have a telescope in space that can directly probe and characterize the formation of those first galaxies.
-This is a key goal of JWST and observational cosmology because the first galaxies probe unknown physics related to star formation (e.g., the efficiency of how gas is converted into stars in early dark matter halos e.g., Behroozi & Silk 2015; Gnedin 2016; Vogelsberger et al.
-2020; Kannan et al.
-2023; Wilkins et al.
-2023) and early structure formation (e.g., Dayal & Ferrara, 2018; Khimey et al., 2021; Lovell et al., 2023; Boylan-Kolchin, 2023).
-While the low metallicity in this first generation of galaxies should slow H formation and decrease star-formation efficiency (e.g., Krumholz & Dekel, 2012), the gas-rich, compact nature of early galaxies may result in increased gas surface densities and star-formation efficiencies (e.g., Krumholz et al., 2009; Ostriker & Shetty, 2011; Somerville et al., 2015).
-There are a wide variety of theoretical predictions for the redshift evolution of the UV luminosity function of galaxies (e.g., Behroozi & Silk, 2015; Mason et al., 2015; Tacchella et al., 2018; Kannan et al., 2023; Wilkins et al., 2023; Katz et al., 2021) These different models typically lead to similar outcomes at regarding number counts and the cosmic star-formation rate density, but they diverge more and more toward higher redshifts due to different treatments of cooling, star formation efficiency, and feedback.
-Similarly, predictions for the formation and evolution of supermassive black holes vary widely (Habouzit et al., 2022).
+This is a key goal of JWST and observational cosmology because the first galaxies probe unknown physics related to star formation (e.g., the efficiency of how gas is converted into stars in early dark matter halos e.g., P.
+S.
+Behroozi & J.
+Silk 2015; N.
+Y.
+Gnedin 2016; M.
+Vogelsberger et al.
+2020; R.
+Kannan et al.
+2023; S.
+M.
+Wilkins et al.
+2023) and early structure formation (e.g., P.
+Dayal & A.
+Ferrara, 2018; D.
+Khimey et al., 2021; C.
+C.
+Lovell et al., 2023; M.
+Boylan-Kolchin, 2023).
+While the low metallicity in this first generation of galaxies should slow H2 formation and decrease star-formation efficiency (e.g., M.
+R.
+Krumholz & A.
+Dekel, 2012), the gas-rich, compact nature of early galaxies may result in increased gas surface densities and star-formation efficiencies (e.g., M.
+R.
+Krumholz et al., 2009; E.
+C.
+Ostriker & R.
+Shetty, 2011; R.
+S.
+Somerville et al., 2015).
+There are a wide variety of pre-flight theoretical predictions for the redshift evolution of the UV luminosity function of galaxies (e.g., P.
+S.
+Behroozi & J.
+Silk, 2015; C.
+A.
+Mason et al., 2015; S.
+Tacchella et al., 2018; R.
+Kannan et al., 2023; S.
+M.
+Wilkins et al., 2023; H.
+Katz et al., 2021) These different models typically lead to similar outcomes at regarding number counts and the cosmic star-formation rate density, but they diverge more and more toward higher redshifts due to different treatments of cooling, star formation efficiency, and feedback.
+Similarly, predictions for the formation and evolution of supermassive black holes vary widely (M.
+Habouzit et al., 2022).
 With JWST, we have the opportunity to push to the redshift frontier of where the physics of these extreme settings cause large differences, factors of 10 and more.
@@
 The exploration of is the most direct way to constrain the onset of cosmic dawn.
-Together with the constraints on the star-formation efficiency and feedback processes at , JWST can bridge the gap to ongoing 21cm experiments—such as EDGES (Bowman et al., 2018), SARAS 3 (Singh et al., 2022) and REACH (de Lera Acedo et al., 2022)—and thereby creating new synergies to constrain the formation of the first stars and galaxies.
-In addition to delivering number density constraints at , JWST observations will constrain the properties of these first galaxies.
+Together with the constraints on the star-formation efficiency and feedback processes at , JWST can bridge the gap to ongoing 21cm experiments—such as EDGES (J.
+D.
+Bowman et al., 2018), SARAS 3 (S.
+Singh et al., 2022) and REACH (E.
+de Lera Acedo et al., 2022)—and thereby creating new synergies to constrain the formation of the first stars and galaxies.
+In addition to delivering number density constraints at , JWST observations constrain the properties of these first galaxies.
 Specifically, these data are sensitive to the total rest-frame UV luminosity, UV continuum slope , and the morphology of those first structures.
-The UV continuum slope is a well-known diagnostic for the stellar populations in galaxies (Bouwens et al., 2009; Wilkins et al., 2011; Finkelstein et al., 2012; Bhatawdekar & Conselice, 2021), primarily affected by the dust attenuation and metallicity.
+The UV continuum slope is a well-known diagnostic for the stellar populations in galaxies (R.
+J.
+Bouwens et al., 2009; S.
+M.
+Wilkins et al., 2011; S.
+L.
+Finkelstein et al., 2012; R.
+Bhatawdekar & C.
+J.
+Conselice, 2021), primarily affected by the dust attenuation and metallicity.
 Together with the UV luminosity, one can derive dust-corrected star-formation rates (SFRs) and infer the cosmic SFR density at .
-Furthermore, an estimate of will probe the chemical enrichment and dust physics in the earliest galaxies, extending studies at from lower redshifts (e.g., Ferrara et al., 2022).
-Very blue UV slopes are a potential signposts of Population III stars or accreting black holes (Schaerer, 2003; Dayal & Ferrara, 2012).
-The structure and morphology of galaxies is a well-known tracer of galaxy assembly (e.g., Conselice, 2014).
-JWST images reveal marginally-resolved galaxies at with compact ( pc) sizes (Robertson et al., 2023; Ono et al., 2023).
-Deep imaging at will deliver rest-UV size constraints down to a resolution of 300 pc.
-Measurements of the size and shape (even upper limits) are a sensitive probe of stellar and black hole feedback in those galaxies (e.g., Pillepich et al., 2019; Wu et al., 2020) and could potentially probe the filamentary structure of the gas accretion (e.g., Pandya et al., 2019; Ceverino et al., 2023).
-Further, one will be sensitive to the multiplicity and close-pair fraction of galaxies at , probing the importance of mergers and star-formation in filaments in these early galaxies (e.g., Wang & Loeb, 2018; O’Leary et al., 2021).
+Furthermore, an estimate of probes the chemical enrichment and dust physics in the earliest galaxies, extending studies at from lower redshifts (e.g., A.
+Ferrara et al., 2022).
+Very blue UV slopes are a potential signposts of Population III stars or accreting black holes (D.
+Schaerer, 2003; P.
+Dayal & A.
+Ferrara, 2012).
+The structure and morphology of galaxies is a well-known tracer of galaxy assembly (e.g., C.
+J.
+Conselice, 2014).
+JWST images reveal marginally-resolved galaxies at with compact ( pc) sizes (B.
+E.
+Robertson et al., 2023; Y.
+Ono et al., 2023).
+Deep imaging at can deliver rest-UV size constraints down to a resolution of 300 pc.
+Measurements of the size and shape (even upper limits) are a sensitive probe of stellar and black hole feedback in those galaxies (e.g., A.
+Pillepich et al., 2019; X.
+Wu et al., 2020) and could potentially probe the filamentary structure of the gas accretion (e.g., V.
+Pandya et al., 2019; D.
+Ceverino et al., 2023).
+Further, one can be sensitive to the multiplicity and close-pair fraction of galaxies at , probing the importance of mergers and star-formation in filaments in these early galaxies (e.g., X.
+Wang & A.
+Loeb, 2018; J.
+A.
+O’Leary et al., 2021).
 3.2 Observational Challenges and a Method to Reach
 There are important reasons why pushing for F200W () dropouts is substantially harder than F115W ( or F150W () dropouts: 1) Such galaxies are fainter and rarer.
-2) The frequency of extreme emission-line galaxies increases with redshift, and JWST clearly has shown a substantial population of them at , where they can create SEDs that differ from those of purely stellar populations, boosting the rest-optical and enhancing the apparent Balmer drop.
+2) The frequency of extreme emission-line galaxies increases with redshift, and JWST clearly has shown a substantial population of them at , where they can create spectral energy distributions (SEDs) that differ from those of purely stellar populations, boosting the rest-optical and enhancing the apparent Balmer drop.
 3) The dichroic gap in NIRCam leaves a wider spacing between F200W and F277W that makes it harder to exclude a mildly reddened Balmer-drop galaxy at , particularly with [OIII] potentially enhancing F277W.
-4) Many Cycle 1 programs include only F277W, F356W, F410M, and F444W.111Notable exceptions are JEMS (Williams et al., 2023) and CANUCS (Withers et al., 2023).
+4) Many Cycle 1 programs include only F277W, F356W, F410M, and F444W.111Notable exceptions are JEMS (C.
+C.
+Williams et al., 2023) and CANUCS (S.
+Withers et al., 2023).
 However, the filters beyond 4 microns have reduced sensitivity due to increased zodiacal foreground, so faint galaxies may only be well-detected in the bluer two bands, which in turn can be filled with [OIII] and H.
@@
 In particular, this search must 1) push very deep in the 2 micron region; 2) increase the spectral resolution to isolate the dropout; and 3) include additional filters in the 2.5–4 micron region to eliminate contamination from mid-redshift emission-line galaxies.
-We believe that this combination of high-SNR detections, well-sampled dropout features, and systematic control over contaminants can reach the high standard of proof that claims of galaxies will require.
-The new imaging will achieve these requirements by combining the usual NIRCam wide-band filters with well-chosen medium-band filters.
+We expect that this combination of high-SNR detections, well-sampled dropout features, and systematic control over contaminants can reach the high standard of proof that claims of galaxies will require.
+The new imaging achieves these requirements by combining the usual NIRCam wide-band filters with well-chosen medium-band filters.
 By observing very deep in F182M, F210M, and F250M, the photometry can isolate a sharp break within a narrower and hence more distinctive region.
-Further, using the combination of medium and wide-bands in the long-wave (LW) channels will reveal strong emission lines that could supply false positive Balmer breaks.
-We illustrate this strategy by considering the candidate in the CEERS field (Donnan et al., 2023), scaled down by a factor of 30 in flux and assuming the noise levels of programs 1210 & 3215.
-While now proved to be at (Arrabal Haro et al., 2023a), this object was a plausible candidate, showing a drop in F200W by a factor of 4 relative to F277W, no detection in F150W, and a blue continuum in the LW filters.
-However, this photometry was also a match to models at with strong optical emission lines and reddening suppression in the UV (Naidu et al., 2022a).
+Further, using the combination of medium and wide-bands in the long-wave (LW) channels reveals strong emission lines that could supply false positive Balmer breaks.
+We illustrate this strategy by considering the candidate in the CEERS field (C.
+T.
+Donnan et al., 2023), scaled down by a factor of 30 in flux and assuming the noise levels of programs 1210 & 3215.
+While now proved to be at (P.
+Arrabal Haro et al., 2023b), this object was a plausible candidate, showing a drop in F200W by a factor of 4 relative to F277W, no detection in F150W, and a blue continuum in the LW filters.
+However, this photometry was also a match to models at with strong optical emission lines and reddening suppression in the UV (R.
+P.
+Naidu et al., 2022b).
 Figure 2 shows how these two hypotheses are starkly differentiated by the NIRCam medium bands: the Lyman break is much steeper than a dust-enhanced Balmer break, and the medium bands easily distinguish continuum from emission lines.
-Shifting the true galaxy in redshift and repeating the fits, Figure 3 demonstrates that this suite of filters provides very effective recovery of candidates at between 13 and 20.
+The fits shown in Figure 2 to the scaled Cycle 1 wide-band photometry use the Prospector galaxy spectrum modeling code (B.
+D.
+Johnson et al., 2021).
+The Prospector models use FSPS (C.
+Conroy et al., 2009) with the MIST isochrones (J.
+Choi et al., 2016), the MILES stellar spectral library (P.
+Sánchez-Blázquez et al., 2006), and nebular emission following N.
+Byler et al.
+(2017).
+The fitted model includes a five-bin flexible SFH with broad priors on the ratio of SFR in adjacent bins following J.
+Leja et al.
+(2019); S.
+Tacchella et al.
+(2022) and normalization of a M.
+Kriek & C.
+Conroy (2013) dust attenuation curve.
+For the nebular emission, we adopt uniform priors on the log of the nebular ionization parameter and log of the gas-phase metallicity, which is allowed to vary independently of the stellar metallicity.
+Attenuation by the IGM follows P.
+Madau (1995).
+The resulting model space can produce a wide range of galaxy SEDs, making it a compelling test of the ability to reject low-redshift interpretations.
+We fit within several relatively narrow redshift ranges and tune the number of live points used for nested sampling with dynesty (J.
+S.
+Speagle, 2020; S.
+Koposov et al., 2022) to help identify narrow local minima in redshift, which can otherwise be missed (e.g.
+R.
+P.
+Naidu et al., 2022b).
+Shifting the true galaxy in redshift and repeating the fits, Figure 3 demonstrates that this suite of filters provides very effective recovery of candidates at between 15 and 20.
+Here we show that for these reference galaxies, a strategy with medium bands produces a much larger rejection of the low-redshift hypothesis.
+This plot uses a fixed normalization of F277W signal-to-noise ratio, but brighter or fainter galaxies would simply scale as the square of the flux.
+What is important is the ratio of the between the two strategies, as this is proportional to how one would have the scale the exposure times in the background-noise limited case.
+One sees that at , this ratio is about 4 times higher, indicating that the medium-band strategy is more performant.
+We note that these galaxy models are based on known stellar populations and do not include the extreme Balmer-break phenomenon associated to black-hole stars (R.
+P.
+Naidu et al., 2025a; A.
+de Graaff et al., 2025), plausibly connected to Little Red Dots (J.
+Matthee et al., 2024).
+Stronger Balmer breaks of course are harder to distinguish from high-redshift candidates, although medium-band photometry will help to distinguish a very abrupt Lyman- break from the smoother decline observed redward of Balmer break in these AGN (R.
+P.
+Naidu et al., 2025a; A.
+de Graaff et al., 2025).
+Such objects are rare and we don’t yet know about their abundance at faint flux levels.
+Importantly, they have strong rest-optical lines that should be readily measurable in spectra.
 3.3 Filters, Exposure Times, and Depths
@@
 | Time (ks) | (ks) | (ks) | (nJy) | redshift | ||
-| F090W | — | — | 55.2 | 55.2 | 3.01 | |
-| F115W | — | — | 72.9 | 72.9 | 2.38 | |
-| F150W | — | — | 55.2 | 55.2 | 2.14 | |
-| F162M | 30 | 82.5 | — | 82.5 | 2.62 | |
+| F090W | — | — | 55.2 | 55.2 | 3.13 | |
+| F115W | — | — | 72.9 | 72.9 | 2.36 | |
+| F150W | — | — | 55.2 | 55.2 | 2.15 | |
+| F162M | 30 | 82.5 | — | 82.5 | 2.80 | |
 | F182M | 60 | 165.0 | — | 165.0 | 1.59 | |
-| F200W | — | — | 38.7 | 38.7 | 2.25 | |
-| F210M | 45 | 123.8 | — | 123.8 | 2.15 | |
-| F250M | 60 | 165.0 | — | 165.0 | 2.40 | |
-| F277W | — | — | 47.0 | 47.0 | 1.90 | |
+| F200W | — | — | 38.7 | 38.7 | 2.24 | |
+| F210M | 45 | 123.8 | — | 123.8 | 2.13 | |
+| F250M | 60 | 165.0 | — | 165.0 | 2.29 | |
+| F277W | — | — | 47.0 | 47.0 | 1.86 | |
 | F300M | 45 | 123.8 | — | 123.8 | 1.80 | |
-| F335M | 30 | 82.5 | 24.8 | 107.3 | 1.81 | |
-| F356W | — | — | 38.7 | 38.7 | 1.95 | |
-| F410M | — | — | 55.2 | 55.2 | 3.30 | |
-| F444W | — | — | 56.4 | 56.4 | 2.65 | |
+| F335M | 30 | 82.5 | 24.8 | 107.3 | 1.84 | |
+| F356W | — | — | 38.7 | 38.7 | 2.05 | |
+| F410M | — | — | 55.2 | 55.2 | 3.56 | |
+| F444W | — | — | 56.4 | 56.4 | 2.95 | |
 Because of the opportunity to re-use the deep imaging of program 1210, we designed program 3215 to return to the same footprint, matching the position angle exactly.
@@
 The driving requirement is to go deep in F182M and F210M, as potential dropout bands.
-In total, F182M, F200W, and F210M will reach 91 hrs of exposure time, far more than the 11 hrs that 1210 & 1180 provide and much deeper than any other Cycle 1 program (to our knowledge, program 2079 being deepest with 17 hrs in F200W).
+In total, F182M, F200W, and F210M reached 91 hrs of exposure time, far more than the 11 hrs that 1210 & 1180 provide and much deeper than any other Cycle 1 program (to our knowledge, program 2079 being deepest with 17 hrs in F200W).
 We split the time roughly evenly between these two medium filters, as we do not know at what redshift a candidate might appear.
@@
 While galaxies could be found anywhere, we site this program in GOODS-S to utilize and build upon the amazing legacy value of this field, while utilizing a position angle that allows ultra-deep MSA spectroscopy on the HUDF at a time of low zodiacal background.
+Of course, these medium-band images also support a wide range of other science cases for intermediate-redshift galaxies, notably emission-line morphologies and SED modeling around the Balmer break, as well as time-domain opportunities from the one-year cadence of the observations.
+The results from this high-redshift search are summarized in §3.7, but we continue next with other aspects of the implementation of program 3215.
 3.4 Coordinated Parallels with NIRSpec
@@
 | GOODS-S Deep/HST | 1210 | 1 | 3 | 250 | 100 | 25 | 25 | 25 | 25 |
-| GOODS-S Year 2 | 3215 | 1 | 5 | 228 | 168 | 42 | 168 | ||
+| GOODS-S Year 2 | 3215 | 1 | 5 | 228 | 134 | 34 | 160 | ||
 Program 3215 follows the path of 1210, using the NIRSpec MSA as the prime instrument and NIRCam as the coordinated parallel.
-This results in exceptionally deep spectroscopy on the HUDF and JADES Deep mosaic, about 104 hrs of total exposure, following on the 55 hrs of program 1210 spectroscopy released in Bunker et al.
-(2023a).
+This results in exceptionally deep spectroscopy on the HUDF and JADES Deep mosaic, about 91 hrs of total exposure, following on the 55 hrs of program 1210 spectroscopy released in A.
+J.
+Bunker et al.
+(2024).
 Following the program goal of exploring the high-redshift universe, we optimize this time toward the study of high-redshift galaxies.
-Unlike program 1210, which was largely targeted from HST ACS optical and WFC3 infrared imaging (Bunker et al., 2023a) save for four early high-redshift candidates from JADES (Robertson et al., 2023; Curtis-Lake et al., 2023), program 3215 will largely utilize NIRCam-selected targets.
+Unlike program 1210, which was largely targeted from HST ACS optical and WFC3 infrared imaging (A.
+J.
+Bunker et al., 2024) save for four early high-redshift candidates from JADES (B.
+E.
+Robertson et al., 2023; E.
+Curtis-Lake et al., 2023), program 3215 utilizes NIRCam-selected targets.
 This allows a great expansion in the number of targets, as the JADES data is comfortably deeper than even the deepest WFC3 HUDF data and much deeper than the flanking CANDELS imaging.
 Bringing ultra-deep spectroscopy to these faint high-redshift galaxies addresses many windows to understanding the physical processes associated with the early galaxy formation.
-We split the spectroscopic time across three dispersing modes—up to 168 ks on source with G395M (far deeper than program 1210), 168 ks with the Prism, and 42 ks with G140M—to serve a variety of key science drivers in studies of these galaxies.
+We split the spectroscopic time across three dispersing modes—up to 134 ks on source with G395M (far deeper than program 1210), 160 ks with the Prism, and 34 ks with G140M—to serve a variety of key science drivers in studies of these galaxies.
 The choice of obtaining ultra-deep spectra in G395M band comes from the requirement of probing the primary rest-optical lines at , with H down to .
@@
 Finally, we seek to characterize the role of neutral hydrogen in and around the galaxies in the reionization epoch (e.g.
-Jones et al., 2023a; Saxena et al., 2023; Witstok et al., 2023) by measuring the Ly emission line.
+G.
+C.
+Jones et al., 2024; A.
+Saxena et al., 2023; J.
+Witstok et al., 2024) by measuring the Ly emission line.
 The PRISM resolution is particularly coarse at the blue end of the spectrum, such that the grating substantially outperforms it for the detection of narrow lines.
@@
 The medium-deep G140M/F070LP spectra are therefore required and adequate to measure the Ly narrow line better than the prism and protect the damping wing analyses.
-The ultradeep spectra delivered by this program will be uniquely suited to pursue several key science cases, as discussed in the following.
-The G395M data will map the early mass-metallicity relation (MZR) at , reaching the unexplored low-mass regime at .
-Indeed, although major progress has been made with JWST in measuring the MZR and the Fundamental Metallicity Relation at z5 (e.g.
-Curti et al., 2023b, a; Nakajima et al., 2023), the low mass regime remains still a largely uncharted territory, except for rare, highly lensed cases (Vanzella et al., 2023) or cases where masses are not yet robustly determined (Maseda et al., 2023).
-With the ultra-deep spectra, we expect to accurately measure the metallicity via the detection of auroral lines, such as [OIII]4363 (Laseter et al., 2023; Sanders et al., 2023a) for tens of galaxies, and via strong line diagnostics (re-calibrated with the auroral-detected spectra) for fainter galaxies, possibly revealing signatures of Population III (Nakajima & Maiolino, 2022; Wang et al., 2022; Maiolino et al., 2023b).
-The spectra are also expected to provide good detections of the CIII]1907,1909 and OIII]1661,1666 transitions, which will be used to determine the C/O abundance ratio, which (due to the different enrichment timescales of carbon and oxygen) is an important clock for the galaxy star formation history (e.g.
-Maiolino & Mannucci, 2019; Arellano-Córdova et al., 2022; Jones et al., 2023b).
-Recent results, based on a few tens of galaxies, have shown that the gas excitation increases steeply at .
-This trend is likely a consequence of higher ionization parameter, lower metallicity, and a harder radiation field, possibly resulting from stellar -enhancement or contribution from X-ray binaries, or even an excess of cosmic rays (Trump et al., 2023; Cameron et al., 2023; Fujimoto et al., 2023b; Katz et al., 2023; Sanders et al., 2023b; Trussler et al., 2022).
-The 3215 deep spectra will explore these trends towards lower mass galaxies, constraining the physics of the ISM with unprecedented accuracy at these redshifts.
-In addition to providing spectroscopic confirmation of galaxies at the redshift frontier, deep prism spectra of distant galaxies have already illustrated the capability of providing tight constraints on stellar population parameters, finding very young ages (a few tens of Myr), very low metallicities (down to ), high yields of ionizing photons, low dust attenuation (Curtis-Lake et al., 2023; Atek et al., 2023).
-Yet, some of these constraints are still highly uncertain and often limited to relatively massive galaxies.
-These ultra-deep spectra will provide much tighter constraints on these parameters and push to much lower masses.
-For about 50 galaxies at z5 with S/N10 per resolution element on the continuum, it will even be possible to derive star formations histories (SFHs Looser et al., 2023a, b), hence testing the highly stochastic behavior of primeval galaxies expected by some models (Ceverino et al., 2018).
-JWST spectroscopy has opened the exploration of the realm of intermediate mass black holes at high redshift, by detecting tens of AGN hosting black holes in the mass range between and , well below the BH mass range probed by quasar surveys (Furtak et al., 2023; Greene et al., 2023; Kocevski et al., 2023; Kokorev et al., 2023; Maiolino et al., 2023a; Matthee et al., 2023b; Übler et al., 2023).
-These studies have detected BHs in AGN that are still relatively luminous ( erg/s), as they were limited by the capability of detecting the nebular lines that can distinguish AGN from star forming galaxies.
-The ultra-deep spectra in program 3215 will allow the detection of even weaker active nuclei in early galaxies.
-Such AGNs can be recognized by their high ionization emission lines (NeV, NeIV, HeII), as well as other strong line diagnostics (Brinchmann, 2023; Nakajima & Maiolino, 2022; Tozzi et al., 2023).
-Further, in unobscured cases, we expect to detect the broad component of the H and H lines in grating spectra, separating this from gas outflows by comparing to the strong [OIII] lines.
-Comparing the inferred black hole masses to the host galaxy stellar mass will shed light on the formation channel of the first generation of black holes, differentiating between scenarios such as direct collapse, merging nuclear clusters, and Population III remnants (Habouzit et al., 2022; Volonteri et al., 2022; Pacucci et al., 2023; Schneider et al., 2023; Trinca et al., 2023).
-Target selection for the program 3215 spectroscopy will be described fully in a later paper.
+The reductions of these ultradeep spectra were included in the JADES Data Release 3 (F.
+D’Eugenio et al., 2025a), along with details of the targeting.
 In brief, we have prioritized high-redshift candidates from JADES, progressively working to lower redshift.
-We include rare special targets at high priority, incorporating AGN indicators, quiescent galaxies at high redshift, sub-mm detections, and strong line emitters extracted from slitless spectroscopy from the FRESCO survey (Oesch et al., 2023).
+We included rare special targets at high priority, incorporating AGN indicators, quiescent galaxies at high redshift, sub-mm detections, and strong line emitters extracted from slitless spectroscopy from the FRESCO survey (P.
+A.
+Oesch et al., 2023).
 As reported in Table 2, we could allocate shutters to 228 unique targets.
 These are slightly fewer targets than allocated in program 1210, because in program 3215 we have given priority in re-observing the same targets in the five pointings, so to maximize the total exposure time for each target.
-Figure 4 shows the distribution of redshift and F444W magnitude for the targets that have been allocated shutters in the NIRSpec observation, illustrating that a significant number of them are very faint ().
-The color coding gives the total exposure time on each target (for the prism configuration), showing that many targets are observed for the full integration time of 168 ksec by being re-allocated shutters in the five pointings.
-We note that the three highest redshift galaxies from Curtis-Lake et al.
+Figure 4 shows the distribution of redshift and F444W magnitude for the targets that were allocated shutters in the NIRSpec observation, illustrating that a significant number of them are very faint ().
+The color coding gives the total exposure time on each target (for the prism configuration), showing that many targets are observed for the full integration time by being re-allocated shutters in the five pointings.
+We note that the three highest redshift galaxies from E.
+Curtis-Lake et al.
 (2023) are targeted for re-observation.
-3.5 Additional Legacy Science
-In addition to the primary science of probing the uncharted regime and revealing galaxy physics with ultra-deep spectroscopy in the parallel observations, the JOF data will provide a rich legacy data set of medium-band imaging and ultra-deep spectroscopy.
-We highlight here some examples of these opportunities.
-The deep 1.6–4.1 m medium-band imaging will allow robust identification of quiescent galaxies via their Balmer breaks down to stellar mass.
-These data will constrain the evolution of the quiescent mass function and measure morphological gradients in Balmer break strengths (Ji et al., 2023), providing crucial constraints on the quenching mechanisms (e.g., Carnall et al., 2023; Nanayakkara et al., 2022; Santini et al., 2022; Marchesini et al., 2023; Looser et al., 2023a, b).
-The exquisite depth will also probe dwarf neighbors of larger mid-redshift galaxies, permitting analysis of their stellar populations and star-formation rates (Suess et al., 2023).
-Turning to star-forming galaxies, by comparing the short-time-scale star-formation rates inferred from the measurement of H emission in the medium-bands to the longer time-scale values inferred from the rest-UV continuum, these data can test whether burstiness increases out to (e.g., Emami et al., 2019; Faisst et al., 2019; Reddy et al., 2023; Looser et al., 2023b).
-Bursts are hypothesized to be a critical driver of the strong emission lines observed at high- (e.g., Stark et al., 2013; Faucher-Giguère, 2018).
-With this very deep imaging, one can also map the line emission and hence star-formation rates across the face of these galaxies.
-The data will also be an opportunity for the infrared time-domain, as the return to the deep Cycle 1 field with very deep year 2 imaging will reveal variability at micron on a timescale of year with medium-band resolution.
-This will probe supernovae (SNe) at (and potentially beyond), for which the SN emission shifts to micron with a peak brightness down to AB mag, as well as detect some AGN by their photometric variability.
 3.5 Implementation
-Program 3215 is implemented in the same manner as 1210, described in Eisenstein et al.
+Program 3215 was implemented in the same manner as 1210, described in D.
+J.
+Eisenstein et al.
 (2023).
@@
 Each visit contains 9 disperser/filter pairs, and each of these is observed for a 3-shutter nod.
-Each exposure contains 2 integrations for NIRSpec (19 groups of NRSIRS2 readout, Rauscher et al.
+Each exposure contains 2 integrations for NIRSpec (19 groups of NRSIRS2 readout, B.
+J.
+Rauscher et al.
 2017, 1400 s each) as well as 2 integrations for NIRCam (7 groups of DEEP8 readout, 1375 s each).
@@
 These are paired with 6 NIRCam exposures of F162M & F335M, 12 exposures of F182M & F250M, and 9 exposures of F210M & F300M.
-The total NIRCam exposure times are summarized in Table 1.
-The exact pointings of the 5 MSA pairs are chosen to place a small number of candidates onto all of the designs, while providing pixel diversity for both NIRCam and NIRSpec.
-We keep the total spacing small, under 2, as this increases the number of targets that can be observed on legal shutters in most of the designs.
-We also use the same position angle as 1210 (V3 PA of 321) and minimize the displacement of the two programs, to ensure high overlap of 14 filters.
-As of this writing, program 3215 is scheduled to be observed in late October 2023, essentially on the one-year anniversary of the Year 1 observations of programs 1210 and 1180 in this field.
-Because program 3215 has no proprietary period, these data sets will go public at nearly the same time.
+The exact pointings of the 5 MSA pairs were chosen to place a small number of candidates onto all of the designs, while providing pixel diversity for both NIRCam and NIRSpec.
+We keep the total spacing small, under 2′′, as this increases the number of targets that can be observed on legal shutters in most of the designs.
+We also use the same position angle as 1210 (V3 PA of 321∘) and minimize the displacement of the two programs, to ensure high overlap of 14 filters.
+3.6 Operational challenges
+The 3215 program was observed October 16–24, 2023, with the data immediately released.
+Most of this data is of the hoped-for quality and has been used in numerous papers (e.g., B.
+Robertson et al., 2024; K.
+N.
+Hainline et al., 2024b; F.
+D’Eugenio et al., 2024; M.
+Curti et al., 2025).
+A minor data quality issue is that unsurprisingly we see a return of the blobs and oil slick artifacts that were seen in program 1210, described in (D.
+J.
+Eisenstein et al., 2023).
+We mask these in our reductions.
+We note that these features also appear (with small translations) in other GOODS-S pointings at this PA, even when the telescope is moved by over an arcminute, but in no pointings we have reduced at other PAs.
+A recurring challenge with very deep NIRCam data is the print-through of background residuals from imperfect flat fields.
+We encountered this for the wide filters and reported in M.
+J.
+Rieke et al.
+(2023) that improved sky flats were needed to control the problem.
+Unsurprisingly, this also affects the medium-band images, and we have found that pushing to the depth of the JOF requires even better flats.
+We note that the dither pattern dictated by the NIRSpec, with 3-step nods and only small transverse steps between visits, appears to exacerbate the meso-scale residuals in the mosaics, even with 15 distinct pointings.
+We will discuss this in a future data release paper, showing that the issue is largely resolved at the depth of the JOF.
+A more serious problem was that the 5th day of observations were impacted by short circuits in the NIRSpec MSA (T.
+D.
+Rawle et al., 2022).
+These badly flood the first 18 exposures (6 nods) of this visit, making useless all of the grating exposures and 3 of the 12 PRISM exposures.
+The remaining 9 PRISM exposures were not affected.
+The total NIRCam exposure times summarized in Table 1 do not include these impacted exposures.
+This problem also created substantial stray-light banding in long-wave module A of the first 6 NIRCam exposures, all F335M.
+The other 12 short-impacted exposures do not show effects, perhaps because the filter changed to F250M, nor is there visible impact in the short-wave data.
+This banding must be masked.
+The imaging observations also encountered substantial persistence in visits 1, 2, and 4.
+In visit 1, the first six exposures, all F162M, have diffuse persistence in modules A3, B3, and B4, typical of a large change in background.
+In visits 2 and 4, the A3 data are affected by persistence from prior wavefront sensing, revealing the pupil image and in visit 4 the ingress and egress trails of the target star.
+These effects are similar to those encountered in JADES (D.
+J.
+Eisenstein et al., 2023).
+As all of the artifacts described above affect only portions of the field, we leave the total exposure times unchanged in Table 1, but note that the effective exposure time on a given location can be mildly shorter.
+We requested and were approved for reobservations of these 6 exposures, both NIRSpec and F162M+F335M imaging, which were planned to occur in October 2024 at the same PA.
+Unfortunately, this plan was rendered impossible due to an ill-timed telescope safing.
+Facing a delay to October 2025, we developed an alternative plan to instead use the remaining time to conduct spectroscopy on the JOF imaging and to use the NIRCam parallel to create a second shallower field with the same filter set.
+This will be described in § 3.8.
+3.7 High-redshift results from the Origins Field
+With the 2023 data, B.
+Robertson et al.
+(2024) analyzed the UV luminosity function using 9 galaxies in the JOF.
+Unfortunately, no galaxies were found, but the upper limits were used to contribute to the inference of the evolution of the UV luminosity function, measuring a drop by a factor of between and 14.
+N.
+J.
+Adams et al.
+(2025) performed a similar analysis in the field of both candidate recovery and the UV luminosity function.
+However, the data set did reveal the astonishing JADES-GS-z14-0 (S.
+Carniani et al., 2024), now spectroscopically confirmed at (S.
+Carniani et al., 2025; S.
+Schouws et al., 2025a, b).
+This galaxy is very bright (45 nJy) and spatially extended, showing that it is not AGN-dominated.
+The JOF would have detected it easily a factor of 10 fainter or out beyond (B.
+Robertson et al., 2024).
+Importantly, the medium-band imaging in the JOF was central to selecting the object as a high-redshift candidate; its proximity to a nearby galaxy was confusing until the medium-band data clearly showed a steep dropout at a mildly different wavelength than the Balmer break of the neighbor.
+Selection was also encouraged by the deep JADES F770W MIRI imaging, which showed a photometric excess reasonably interpreted as the H-[OIII]5007Å emission-line complex (J.
+M.
+Helton et al., 2025).
+With JADES-GS-z14-0 and the recent MoM-z14 (R.
+P.
+Naidu et al., 2025b), only twice fainter, it is clear that there is a detectable population of galaxies at .
+Indeed, the brightness of these two known galaxies might recommend a wider and shallower strategy.
+But whether bright or faint, the effects of the dichoic wavelength and the presence of strong emission lines in the LW filters, discussed in § 3.2, should be folded into the design and will likely continue to favor the use of medium bands.
+The 3215 imaging data reduction and catalogs will be included in the upcoming JADES data release.
+3.8 Dense-shutter spectroscopy on the JOF
+With the opportunity to replan the remaining time into a window in December 2024, we proposed and were approved to change the pointing so as to conduct multi-object spectroscopy on the JOF.
+At the same time, the NIRCam parallels at this position angle (V3_PA 39∘) fell on deep wide-band imaging from JADES program 1286.
+By doing the same medium-band filter set, we could create a second field with the same 15 filters as the JOF, albeit not as deep.
+This was conducted on December 17–18, 2024.
+This final pointing consists of a single MSA configuration, which was then observed for 9 nods of a 3-high shutter, providing 3 distinct dither locations.
+8 of the nods use long exposures, the same as the original 3215 plan; the 9th is much shorter, only a single integration per position with 6 groups of NIRCam MEDIUM8 readout and 10 groups of NIRSpec NRSIRS2 readout, in order to use the available time allocation.
+The NIRCam imaging provides 3 long nods of F250M, F300M, and F210M, 2 long nods of F162M, 2 long and 1 short nod of F182M, and 1 long and 1 short nod of F335M, this last filter being kept short because JADES also provides some of this filter.
+We also began the visit with 1 long nod of F150W and F356W, arguing to duplicate this filter from program 1286 as a guard against time-variable sources, which we have found to present false positives for dropout galaxy searches in multi-epoch surveys (C.
+DeCoursey et al., 2025).
+For the spectroscopy, we decided to attempt a more experimental use of the NIRSpec MSA.
+JADES had been routinely using designs that create overlapping grating spectra and non-overlapping prism spectra, relying on the prism spectra to reliably sort out which emission lines in the grating data belong to which target.
+Here, we opted to omit the prism spectra and observe a single MSA configuration with many overlapping grating spectra, relying on astrometric variations along the slit and priors from the photometric redshifts to associate the lines.
+In doing so, we achieve very high multiplex, 850 targets on this one configuration, with a focus solely on emission-line science.
+We open shutters on any target that is sufficiently well centered.
+The targets were galaxies with photometric redshifts above 3, whose strong emission lines would fall in the G235M and G395M grating, with flux limits down to 7 in the very deep LW imaging.
+We omit any galaxy brighter than 300 nJy, as those dispersed continua would be bright enough to add noise to the overlapping extractions.
+The approach is motivated by several important technical facts.
+First, the NIRSpec grating spectra are detector-noise limited; it requires over 10 overlapping spectra to generate sufficient background to overcome this.
+This implies that overlapping the spectra does not generate a harmful amount of photon noise.
+Second, the astrometric knowledge from NIRCam gives a strong prior on where lines can occur along the 1.6′′ length of a 3-shutter slit.
+In the wavelength direction, one has additional ways to determine associations from detection of multiple lines and from priors from photometric redshifts and even the detection of the line in the multi-band imaging.
+This association problem is familiar from slitless spectroscopy, such as with the NIRCam grism; indeed, it is easier here because one know which targets were given a shutter.
+Third, while the density of faint galaxies is high enough that continuum spectra are overlapped, the density of detectable emission lines is low enough that one is very far from the confusion limit.
+By using G235M and G395M together on one MSA configuration, we acquire 1.7–5.3 micron spectra, with only small wavelength gaps due to the chip gap.
+We observed 4 long nods in each of G235M and G395M, and a short nod in G395M, totaling 33.6 and 35.8 ksec, respectively.
+The spectral performance was very good, detecting lines down to 2– ergs/cm2/s, and we netted 539 successful redshifts in this one MSA configuration.
+Further description of the method, along with target selection, resulting performance, and the redshift catalog will be presented in D’Eugenio et al.
+(in prep.).
 4 Cycle 3: NIRCam Grism Spectroscopy with Program 4540
 4.1 Science Motivation
-While the investment of JWST imaging time in the JOF is tremendous, this field does not yet have the spectroscopic resources of the HUDF.
-The JOF lies in the flanking depth region of GOODS-S and CANDELS, rather than in the HUDF itself.
-GOODS-S has had substantial spectroscopic attention from ground-based observatories and the 3D-HST survey, but it does not have the ultra-deep data that the HUDF, particularly from the deep VLT/MUSE program.
-The JOF also has only limited overlap with the JWST FRESCO program.
-In December 2023, this area will be targeted with 4-5 NIRSpec MSA configurations from program 1286 and 1287, which will provide coverage of some high-priority objects.
-The NIRCam grism has been shown to produce a compelling capability to measure strong [OIII] and H emission lines from galaxies at (Sun et al., 2023a; Kashino et al., 2023; Matthee et al., 2023a; Oesch et al., 2023; Wang et al., 2023c).
-Such observations would be an excellent pairing to the imaging, providing spectroscopic redshifts, line diagnostics, and dynamical measurements to this deep window on the universe.
-The NIRCam instrument team therefore has decided to devote 32 hours of Cycle 3 guaranteed time toward such a program (program ID 4540, PI: Eisenstein).
-There are a number of key motivations from the science opportunity from these data. Some of these are generic to NIRCam grism data, but many are specific to the opportunities of this special field.
-The core goal is to make make more complete use of the opportunity at that the [OIII] and H lines are often extremely strong.
-Observing only one filter leaves a redshift gap, but observing in both the F356W and F444W filters provides complete coverage at with these two strong lines.
-Further, at –6.6, the data provide access to both lines, which is useful for studying the metallicity of the galaxy populations and which nails the line identification.
-At –6.6, these data will include both H and H, the ratio of which allows us to infer the amount of dust reddening, indeed free from the complicated path-loss corrections that result from the NIRSpec MSA.
-At –5.2, we will measure H in the grism but will also measure [OIII] and H in the F250M, F300M, and F277W imaging.
-This will allow us to extend the measurement of the ratio of [OIII] to Balmer lines to a larger population of galaxies.
-These same strong lines are usually detected in the medium-band imaging, where comparison between the filters allows one to measure the morphology in the line, as distinct from the rest-optical continuum.
-When combined with the grism data at a spectral resolution , one can break the degeneracy between the velocity field and the morphology, providing spatially resolved velocity measurements in some galaxies (Nelson et al., 2023).
-Extending this opportunity, we note that a portion of these program 4540 observations overlap cycle 1 F444W grism observations from FRESCO (Oesch et al., 2023) but with a dispersion direction that differs by about 50.
-This will improve the velocity field measurements, as slitless spectroscopy always has a degeneracy when there is a velocity gradient across the object in the dispersion direction.
-It also provides another way to confirm the association of grism lines to the galaxies in the undispersed images.
-The high spectral resolution of the NIRCam grism yields important dynamical information from the line widths of the emission lines.
-The widest lines can indicate galactic outflows or AGN broad lines, giving critical clues to galactic feedback and the role of black holes in galaxy growth (Matthee et al., 2023b; Maiolino et al., 2023a; Greene et al., 2023; Furtak et al., 2023; Harikane et al., 2023b; Kokorev et al., 2023).
-The spectroscopy also creates important opportunities for the modeling of the detailed infrared SEDs.
-For , these spectroscopically detected galaxies will have their Balmer breaks in the progression of medium bands from program 3215; this will allow careful characterization of star formation histories.
-The detailed SED shapes around the Balmer and 4000Å breaks are key for detecting stars older than about 50 million years, but the effects are subtle and can be confused by degeneracies between the rest-frame colors and the photometric redshifts.
-Limits on the strong Balmer lines also allow one to clean the effects of the weaker Balmer lines around 4000Å from the continuum photometry.
-Finally, the strong lines themselves can be included in the galaxy spectral synthesis analyses, as a constraint on the most recent star formation.
-Some of these galaxies will have very deep MIRI F770W data from program 1180, probing rest-frame 1–2 at these redshifts.
-This will provide more robust constraints on stellar populations and in particular on the stellar mass of these emission-line galaxies.
-We stress the opportunity to compare these stellar masses to the dynamical masses from resolved velocity profiles.
-Although the MIRI data will only detect a subset of the galaxies, that subset can be used to refine the analysis in the NIRCam bands to improve mass determinations overall.
-Moving beyond the modeling of single galaxies, an exciting aspect of the NIRCam grism is the ability to detect a rich pattern of galaxy overdensities at these high redshifts.
-The FRESCO, EIGER, and ASPIRE data sets have shown beautiful results in this regard (Kashino et al., 2023; Witstok et al., 2023; Helton et al., 2023; Wang et al., 2023c; Herard-Demanche et al., 2023; Sun et al., 2023b).
-The patterns are much easier to find in three-dimensional mapping, and the ability of the grism to observe nearly every galaxy in an unbiased way has been crucial for getting to dense enough coverage to show these structures.
-In doing so, we can place the galaxies in their environmental context, connecting all of the imaging information about morphology, SEDs, and emission-line properties to environment.
-We judged this opportunity to measure this large-scale structure context so compelling that we opted to spread this grism data to cover the inter-module gap, even though programs 1210 and 3215 do not cover it, so that we can measure the environment of galaxies that fall near the edge of the imaging modules.
-Indeed, Helton et al.
-(2023) presents the discovery of a substantial arcminute-scale galaxy overdensity with 43 emission-line galaxy members extracted from the south-west corner of the FRESCO data.
-Some of the galaxies fall exactly in the program 4540 footprint, which will extend the mapping beyond the FRESCO boundary.
-Our deeper data will better measure the dynamics of this forming structure through the [OIII] and H detections of these galaxies.
-Finally, while the above focuses on the strong [OIII] and H lines, these grism data will net spectroscopic detections at lower redshift of the weaker rest-infrared lines, confirming redshifts to strong star-forming dwarfs and measuring lines whose equivalent widths are likely too small to claim from the imaging.
-We particularly note the ability to detect Paschen to (and Paschen to ), the comparison of which to the rest-UV continuum measured in the HST imaging in this field and H available in the F150W and F162M (and out to F250M, for P) filters can more directly reveal dust-obscured star formation.
-He I 10830Å and [S III] 9532Å lines can also be detected in vigorous cases, with line identification coming from the medium-band detections of the stronger lines at bluer wavelengths.
+While the investment of JWST imaging time in the JOF is tremendous, this field does not yet have the deep spectroscopic attention that the HUDF has received.
+The NIRCam instrument team therefore decided to devote 32 hours of Cycle 3 guaranteed time to a NIRCam slitless spectroscopy program (program ID 4540, PI: Eisenstein), augmenting the MSA coverage from programs 1286, 1287, and 3215.
+The NIRCam grism has been shown to produce a compelling capability to measure strong [OIII] and H emission lines from galaxies at (F.
+Sun et al., 2023; D.
+Kashino et al., 2023; J.
+Matthee et al., 2023; P.
+A.
+Oesch et al., 2023; F.
+Wang et al., 2023).
+By observing in the F322W2, F356W, and F444W filters, program 4540 produced slitless spectroscopy from 2.4 to 5 m over most of the JOF, covering the [OIII] or H lines at .
+Of course, rest-infrared lines are also detected from brighter low-redshift galaxies as well.
+Early reductions indicate that these data have measured over 1400 redshifts across 0–9, nearly 100 per square arcminute.
+Slitless spectroscopy is an exciting pairing to the JOF deep medium-band imaging, as it provides precise redshifts and line shape measurements to the lines inferred from the imaging.
+Precise redshifts in turn allow the medium-band photometry to make more accurate inferences about the continuum breaks, free from uncertainties as to where the weaker lines might fall.
+Indeed, the measurement of the line emission in the imaging breaks the degeneracy in slitless spectroscopy between astrometric position and wavelength, allowing the combination to measure velocity gradients within a galaxy.
+Further, line shapes are a powerful tool to identify Type I AGN and strong galactic outflows (J.
+Matthee et al., 2024; R.
+Maiolino et al., 2024; J.
+E.
+Greene et al., 2024; L.
+J.
+Furtak et al., 2024; Y.
+Harikane et al., 2023b; V.
+Kokorev et al., 2023).
+A general census of strong line emitters also provides measures of galactic environment (D.
+Kashino et al., 2023; J.
+Witstok et al., 2024; J.
+M.
+Helton et al., 2024; F.
+Wang et al., 2023; T.
+Herard-Demanche et al., 2025; F.
+Sun et al., 2024) for studies of SED properties.
+Indeed, although this program was originally planned only to include F356W and F444W grism data, the presence of an obvious overdensity of F300M-excess sources in the JOF due to two overdensities at and 3.70 encouraged us to split the F356W time to include the F322W2 filter, so that we would have H coverage of these structures.
+In addition, J.
+M.
+Helton et al.
+(2024) presents the discovery of a substantial arcminute-scale galaxy overdensity with 43 emission-line galaxy members extracted from the south-west corner of the GOODS-S FRESCO data (P.
+A.
+Oesch et al., 2023).
+Combining with the Cycle-3 pure-parallel NIRCam grism survey SAPPHIRES (F.
+Sun et al., 2025; F.
+D’Eugenio et al., 2025b), Program 4540 has extended this mapping spatially and add [OIII] detections to the H coverage from FRESCO.
+Further aspects of the data reduction of this grism data and resulting spectral catalog will be presented in Sun et al. (in prep.).
 4.2 Short-wave Data
 We use the opportunity of program 4540 to add deep F070W imaging to the JOF, the final NIRCam broad-band filter and 15th NIRCam filter overall.
-Given the mosaic pattern described below, this will reach exposure times of 16–68 ks, depending on overlaps, mildly shallower than but comparable to the 55 and 73 ks in F090W and F115W, respectively.
-This will extend the JWST coverage of the SEDs and produce optical imaging much deeper than the ACS imaging from GOODS and CANDELS.
-These F070W data provide key enhancements for the core science of the program, as it will provide direct sensitivity to a fainter population of Ly break galaxies at –6.
-This will provide a UV-selected population to be compared to the emission-line selected population in [OIII] and H and to the characterization of Balmer/4000Å breaks from the medium-band data.
+Given the mosaic pattern described below, this reaches exposure times of 16–68 ks, depending on overlaps, mildly shallower than but comparable to the 55 and 73 ks in F090W and F115W, respectively.
+This extends the JWST coverage of the SEDs and produces optical imaging much deeper than the ACS imaging from GOODS and CANDELS.
+These F070W data provide key enhancements for the core science of the program, as it provides direct sensitivity to a fainter population of Ly break galaxies at –6.
+This yields a UV-selected population to be compared to the emission-line selected population in [OIII] and H and to the characterization of Balmer/4000Å breaks from the medium-band data.
 4.3 Implementation
-Driven by the desire to cover most of the 3.1 to 5 m range over the full program 1210 footprint as well as to cover the intermodule gap and a small boundary for environment measurements, we design program 4540 for a tight 22 mosaic pattern as shown in Figure 5.
-The same data will give partial wavelength coverage in a mildly broader area; this bounding area is largely covered by JADES medium-depth NIRCam imaging.
-As redshifts in this field are so valuable, we aim to achieve a depth that is notably fainter than the FRESCO observations.
-In the overlap regions of the 4 tight pointings, we will reach a depth of over 9 hrs of integration time in each filter, compared to 2 hrs in FRESCO F444W.
-The F356W filter data is yet more sensitive because of lower zodiacal background.
-We expect to measure several hundred redshifts at , in addition to the lower redshift galaxies detected in lower luminosity redder lines.
+Driven by the desire to cover the 3 to 5 m range over the full program 1210 footprint as well as to cover the intermodule gap and a small boundary for environment measurements, we designed program 4540 for a tight 22 mosaic.
+As shown in Figure 5, the same data give partial wavelength coverage across 2.4–5.0 m in a mildly broader area; this bounding area is largely covered by JADES medium-depth NIRCam imaging.
+As redshifts in this field are so valuable, we aim to achieve a depth that is notably fainter than the FRESCO observations ( depth of unresolved emission line is erg s-1 cm-2 around 4.3 µm).
+In the overlap regions of the 4 tight pointings, we reach a maximum depth of hrs of integration time in the F444W filter, compared to 2 hrs in FRESCO F444W.
+We had originally planned for an equal amount of time in F356W, but later decided to split those exposures evenly between the F356W and F322W2 filters, noting that the F356W filter data is yet more sensitive because of lower zodiacal background.
+Preliminary analyses suggest that the data have reached a depth for unresolved emission lines of erg s-1 cm-2 around 4.3 µm (with the F444W filter at a median exposure time of 5.4 hrs); erg s-1 cm-2 around 3.6 µm (with the F356W filter alone at a median exposure time of 3.0 hrs); and erg s-1 cm-2 around 2.8 µm (with the F322W2 filter at a median exposure time of 3.0 hrs).
 The grism data are undeflected at 3.95 , so the regions of complete F356W spectra and complete F444W spectra fall in different portions of the NIRCam field of view.
 With the column-direction grism (GRISMC), we use 2 rows of the mosaic stepped so that we can get complete coverage over the full JOF NIRCam imaging footprint.
-We then offset between F356W and F444W mosaics by so that we get complete 3.1 to 5 coverage over a consistent footprint.
+One observation observed in all 3 filters; this produces coverage maps that are displaced as a function of wavelength.
+The other 2 observations were offset, with F444W toward the northwest and F356W/F322W2 toward the southeast, so that we get complete 3 to 5 coverage over the JOF.
 GRISMC is preferred to GRISMR for this purpose because the dispersion direction is the same in the two modules.
-Following the pattern well-used in our earlier GTO program, we are using the DEEP8 readout mode to minimize data volume.
-We will use long integrations (6 or 7 groups, 1159–1375 sec) to minimize overheads and reduce the impact of detector noise.
-We use 6–8 dither positions in each of the mosaic positions to produce good pixel redundancy and achieve the desired exposure time The INTRAMODULEBOX pattern is tightest and most optimal for our purpose.
-Because of allocation constraints, we have slightly different exposure times per pointing, and we use longer exposures on the two F356W grism images with more coverage of the JOF.
+Blueward of 3.1 microns, the F322W2 data (which use the same pointing as F356W) sample an even smaller field, with gaps opening in the footprint at the very blue end.
+We use 6 dither pointings in each of the mosaic pointings, using the INTRAMODULEBOX pattern.
+Because of allocation constraints, we have different exposure times.
+Two of the pointings use DEEP8 readout, with 5 groups for F322W2 and F356W and 8 groups for F444W.
+The other two pointings use 8 groups of MEDIUM8 readout, with F444W being observed twice.
 Although slitless spectroscopy often uses multiple dispersion directions to robustly connect lines to the zero-order location and hence get unique wavelengths, we do not do this here.
-With the sharp NIRCam imaging and accurate grism spectral tracing models, the associations of lines to galaxies detected in the imaging is often unconfused.
-More importantly, this field has exceptionally deep multi-band imaging, and these high-equivalent emission lines (essentially all of the cases) show up easily in the medium-bands from programs 1180, 1210, and 3215.
+With the sharp NIRCam imaging and accurate grism spectral tracing models, the associations of lines to galaxies detected in the imaging is unconfused in most cases.
+More important, because of the deep multi-band imaging, the high-equivalent emission lines (essentially all of the cases) show up easily in the medium-bands from programs 1180, 1210, and 3215.
 This sharply reduces concerns about the ability to robustly match up the line detections to the undispersed images.
+Combining with the emission-line maps from imaging can even break the degeneracy between morphology and velocity fields inherent to slitless spectroscopy (e.g., E.
+Nelson et al., 2024; A.
+L.
+Danhaive et al., 2025).
 Further, the NIRCam imaging yields very high quality photometric redshifts, which reduces the pressure to have a second line detection to produce an identification of the line and hence a redshift.
+That said, the lack of multiple dispersion angles does sharply limit the ability to robustly extract continuum spectra from the grism data.
 Direct images are required by the slitless spectroscopy observing template.
-We use short exposures (5–6 groups of SHALLOW4, 257–311 sec) because this field already has far deeper data.
-We use the same LW filters, but change in SW to F150W and F200W to increase the signal-to-noise in the short exposures and to assure some infrared coverage in the outer portions of field where we may fall off of previous NIRCam imaging.
-In detail, we found that layout of this program in a rigid mosaic was impossible given a hole in guide star coverage.
-Because of the two out-of-field direct images per pointing and the offset between the pointings of the two filters, there are a total of 24 separate visits.
-To achieve schedulability at a consistent position angle, we had to decouple the 8 grism positions and make small (10-20) level adjustments to find guide stars.
-Hence, the mosaic is an irregular grid.
+We use short exposures (6 groups of SHALLOW4, 311 sec) because this field already has far deeper data.
+We use the same LW filters, but change in SW to F150W and F200W to increase the signal-to-noise in the short exposures, to assure some infrared coverage in the southwest edge of field where we may fall in gaps of other NIRCam imaging, and to get a 3rd separate year of data for time-domain studies.
 4.4 Coordinated Parallels with MIRI
-Program 4540 will also produce coordinated parallels for MIRI imaging in the F770W and F1500W filters.
-For each filter, these will reach a depth of about 2 hrs in each of 4 overlapping pointings, with a peak exposure time around 9 hrs.
-We choose to use F770W in the outermost 4 pointings, yielding a contiguous field of about by , while we use F1500W in the inner 4 pointings, yielding a field of by that is contained within the F770W mosaic.
-As with JADES, we use SLOWR1 readout mode to limit the data volume, and we use multiple integrations per NIRCam integration to mitigate saturation.
-These F770W data will extend the study of galaxies and particularly their stellar mass utilizing rest-frame 1–2 measurements, while the the F1500W data will provide insight into the role of AGN in –5 galaxies, as the hot dust in the nuclei can outshine the Rayleigh-Jeans tail of the stellar population.
+Program 4540 also produced coordinated parallels for MIRI imaging in five filters.
+The longer grism observation with all 3 LW filters is used for F1500W and F2100W; this produces the tightest footprint about 2′ square with total depth of about 7 ks per filter .
+The other two observations use F770W and F1000W, producing a footprint centered on the first with total depth of about 5 ks per filter.
+We note that one of the four MIRI F1000W pointings was lost because of insufficient disk space on the spacecraft, resulting in reduced coverage and depth at the southeastern corner of the mosaics.
+All of the direct imaging use F560W, producing a mildly larger footprint; the increased sensitivity of F560W will compensate for the shallower depth of 1–2 ks.
+The footprints of coordinated MIRI parallels are also shown in Figure 5.
+The longer observation uses FASTR1 readout; the other two use SLOWR1 to limit the data volume.
+Aside from F560W, we use multiple integrations per NIRCam integration to mitigate saturation.
+In detail, the number of integrations and groups are 3 & 11, 3 & 11, 3 & 71, and 14 & 31 in F770W, F1000W, F1500W, and F2100W, respectively; F560W uses either 13 slow or 100 fast groups.
+The JWST ETC forecasts 5-sigma depths of 25.0-25.3, 25.5, 25.0, 24.5, and 22.9 AB mag in the F560W, F770W, F1000W, F1500W, and F2100W bands, respectively, assuming a point source with an 0.6′′ diameter aperture.
+These F770W and F1000W data will extend the study of galaxies and particularly their stellar mass utilizing rest-frame 1–2 measurements, while the longer wavelength data will provide insight into the role of AGN in –5 galaxies, as the hot dust in the nuclei can outshine the Rayleigh-Jeans tail of the stellar population.
 Finding such sources may be the most powerful means to identify heavily obscured AGN at .
 However, the location of this parallel field does fall into an area currently lacking in NIRCam and HST imaging.
-Of course, the extended CDF-S does have deep ground-based imaging (such as the deep 5-year stack from the Dark Energy Survey SNe program) as well as imaging from Spitzer, which can support near-term use.
+Of course, the extended CDF-S does have deep ground-based imaging (such as the deep 5-year stack from the Dark Energy Survey SNe program and the recently released Rubin DP2) as well as imaging from Spitzer, which can support near-term use.
 We expect that this field is likely to be imaged with NIRCam in the future, as its location is easy to reach as a parallel when NIRSpec is pointed at the exceptional JOF NIRCam imaging, indeed at the position angle such that the NIRSpec quadrants align to the NIRCam geometry.
@@
 In this paper, we present the second imaging data release of JADES, including the GOODS-S imaging observed in September and October 2022.
-The first release (Rieke et al., 2023) included the NIRCam imaging of the Deep Prime program on the HUDF.
+The first release (M.
+J.
+Rieke et al., 2023) included the NIRCam imaging of the Deep Prime program on the HUDF.
 This second release includes the deep NIRCam parallel from program 1210 and the medium-depth mosaic that flanks these two imaging fields.
-Caveats regarding data quality and some missing data due to short circuits in the NIRSpec MSA and one skipped visit are described in Eisenstein et al.
+Caveats regarding data quality and some missing data due to short circuits in the NIRSpec MSA and one skipped visit are described in D.
+J.
+Eisenstein et al.
 (2023).
 Figure 6 shows a three-color image of the field.
-The total area of the JADES imaging (excluding the FRESCO-only region) is 67.7 arcmin, and the catalogs contain 94,000 distinct objects.
+The total area of the JADES imaging (excluding the FRESCO-only region) is 67.7 arcmin2, and the catalogs contain 94,000 distinct objects.
 A summary of the image reduction and catalog processing follows.
-To create calibrated imaging and object photometry catalogs, the NIRCam data are reduced in the same manner as described in Rieke et al.
+To create calibrated imaging and object photometry catalogs, the NIRCam data are reduced in the same manner as described in M.
+J.
+Rieke et al.
 (2023, hereafter R23) with a few modifications.
@@
 The JWST calibration reference data file used during processing was jwst_1130.pmap, which adjusted the photometric zeropoints slightly relative to the first JADES data release.
-We have checked that the flux ratios of objects shared between the current data release and Rieke et al.
+We have checked that the flux ratios of objects shared between the current data release and M.
+J.
+Rieke et al.
 (2023) are broadly consistent with the changes in photometric zeropoint.
@@
 These show the deep parallel field of program 1210, along with the year 1 data in the Deep and Medium mosaics of program 1180.
-Year 2 observations will double the depth of the Deep mosaic, add flanking Medium depth around the Deep region, and extend the footprint with parallel NIRCam imaging from program 1286 and 1287 (Eisenstein et al., 2023).
+Year 2 observations doubled the depth of the Deep mosaic, added flanking Medium depth around the Deep region, and extended the footprint with parallel NIRCam imaging from program 1286 and 1287 (D.
+J.
+Eisenstein et al., 2023).
 Figure 8 shows maps of the corresponding 5 sensitivity for F115W, F200W, and F356W in diameter apertures, with aperture corrections applied.
 These maps are computed by placing randomized apertures throughout the fully reduced mosaic images, after masking sources and surrounding pixels with significant flux.
-As reported by Rieke et al.
+Such errors are preferred to formal errors derived from pixel-level uncertainties because of the correlations between pixels induced by the mosaicking.
+As reported by M.
+J.
+Rieke et al.
 (2023), the 1180 Deep region reaches F200W depth of mag AB.
-To enhance the multi-band space-based photometry available for sources, we also perform photometry on the Hubble Legacy Field HST/ACS F435W, F606W, F775W, and F814W mosaics and HST/WFC3 F105W, F125W, F140W, and F160W mosaics (Illingworth et al., 2016) after rectifying them to our astrometric frame.
-We reprocess and release JWST/NIRCam F182M and F210M mosaics from data acquired by the JWST Extragalactic Medium-band Survey (JEMS; Williams et al., 2023) and First Reionization Spectroscopically Complete Observations (FRESCO; Oesch et al., 2023) programs.
+To enhance the multi-band space-based photometry available for sources, we also perform photometry on the Hubble Legacy Field HST/ACS F435W, F606W, F775W, and F814W mosaics and HST/WFC3 F105W, F125W, F140W, and F160W mosaics (G.
+Illingworth et al., 2016; K.
+E.
+Whitaker et al., 2019) after rectifying them to our astrometric frame.
+We reprocess and release JWST/NIRCam F182M and F210M mosaics from data acquired by the JWST Extragalactic Medium-band Survey (JEMS; C.
+C.
+Williams et al., 2023) and First Reionization Spectroscopically Complete Observations (FRESCO; P.
+A.
+Oesch et al., 2023) programs.
 We additionally reprocess and release JWST/NIRCam F430M, F460M, and F480M images from data acquired by JEMS and combine reprocessed JWST/NIRCam F444W exposures originally acquired by FRESCO with our F444W stack.
 We include photometry from all JWST filters in our released catalogs, where available.
-The construction of the photometric catalogs largely follows that described by Rieke et al.
+The construction of the photometric catalogs largely follows that described by M.
+J.
+Rieke et al.
 (2023).
@@
 Catalogs are constructed from the resulting segmentations, with automated and manual masking of stars and diffraction spikes applied.
-SExtractor (Bertin & Arnouts, 1996) windowed positions are used as object centroids.
-Aperture photometry is measured using a range of circular apertures, two Kron apertures, and segmentations as discussed in Rieke et al.
+SExtractor (E.
+Bertin & S.
+Arnouts, 1996) windowed positions are used as object centroids.
+Aperture photometry is measured using a range of circular apertures, two Kron apertures, and segmentations as discussed in M.
+J.
+Rieke et al.
 (2023).
 Sky background uncertainties are computed by placing random apertures throughout the images and directly from the jwst pipeline ERR image layers when available.
-Aperture corrections are computed from model point-spread-functions constructed following Ji et al.
+Aperture corrections are computed from model point-spread-functions constructed following Z.
+Ji et al.
 (2024).
 Flags are set to indicate bad pixels that may impact object segments, whether sources may be influenced by brighter neighbors, and whether stars identified in the image may impact them or if the objects themselves are consistent with stars.
-In addition to the photometry, we have computed photometric redshifts following the methods described in Rieke et al.
-(2023) and Hainline et al.
-(2023).
-In particular, these use the EAZY code (Brammer et al., 2008) with templates as described in Hainline et al.
-(2023) and incorporating percent-level adjustments to the photometric zeropoints based on the global median offset of each filter in a first-round of fitting.
+In addition to the photometry, we have computed photometric redshifts following the methods described in M.
+J.
+Rieke et al.
+(2023) and K.
+N.
+Hainline et al.
+(2024a).
+In particular, these use the EAZY code (G.
+B.
+Brammer et al., 2008) with templates as described in K.
+N.
+Hainline et al.
+(2024a) and incorporating percent-level adjustments to the photometric zeropoints based on the global median offset of each filter in a first-round of fitting.
 For convenience, the photometric redshift catalog is included as FITS extension of the photometry catalog.
+A comparison to spectroscopic calibrators was plotted for DR1 in M.
+J.
+Rieke et al.
+(2023) and K.
+N.
+Hainline et al.
+(2024a), showing good performance in the aggregate.
+Of course, calibrations for faint populations are meager, and one should be careful about how outliers, particularly of unusual colors, might impact one’s results.
+We note that with the medium-band photometry, galaxies with strong rest-optical emission lines often have more robust photometric redshifts, so performance is not a simple function of continuum magnitude.
 The images and catalogs are available on the Mikulski Archive for Space Telescopes (MAST) at https://archive.stsci.edu/hlsp/jades, with DOI 10.17909/z2gw-mk31.
-We will incorporate the 3215 imaging using this same pipeline in an incremental release.
+We will incorporate the 3215 imaging using this same pipeline in an upcoming release.
 6 Discussion
-We have described the executed and approved plans for JWST observations of the JADES Origins Field, spanning the first three mission cycles.
-In total, the time to be invested in this single NIRCam footprint of 9 arcmin is already very large, roughly 360 open-shutter hours of exposure time across in 15 NIRCam imaging filters, 20 hours of NIRCam slitless spectroscopy, and over 170 hours of MIRI imaging.
+We have described the JWST observations of the JADES Origins Field, spanning the first three mission cycles.
+In total, the time invested in this single NIRCam footprint of 9 arcmin2 is already very large, roughly 360 open-shutter hours of exposure time across in 15 NIRCam imaging filters, 20 hours of NIRCam slitless spectroscopy, and over 170 hours of MIRI imaging.
 This is the most intensively observed JWST field yet, designed to probe the redshift frontier beyond .
-Figure 9 shows an small portion of the JOF imaging data from program 1210 to highlight the opportunities of the deep NIRCam medium-band data, here F335M and F410M.
+Figure 9 shows a small portion of the JOF imaging data from program 1210 to highlight the opportunities of the deep NIRCam medium-band data, here F335M and F410M.
 Strong emission lines are revealed easily in these images, with beautiful morphological detail.
-This reach will be greatly extended in program 3215 with 5 new medium bands spanning 1.6 to 3.0 .
+This reach was greatly extended in program 3215 with 5 new medium bands spanning 1.6 to 3.0 .
 The location of the JOF in close proximity to the HUDF and the Deep Prime JADES NIRCam imaging mosaic adds considerable value to the GOODS-S region.
@@
 By construction, further NIRSpec observations of the HUDF can place further NIRCam exposure time in the JOF region; the position angle is at a period of low zodiacal background and is allowable with the micrometeorite constraint.
-Indeed, program 3215 is using this opportunity to conduct exceptionally deep spectroscopy (104 hrs of total exposure) in the HUDF, focusing on the physics of high-redshift galaxies.
+Indeed, program 3215 used this opportunity to conduct exceptionally deep spectroscopy (91 hrs of total exposure) in the HUDF, focusing on the physics of high-redshift galaxies.
 Further, observations with NIRISS or MIRI on the JOF are easily paired with NIRCam observations on and around the HUDF, within the JADES Deep NIRCam mosaic.
@@
 These data are available at https://archive.stsci.edu/hlsp/jades and a feature-rich FITSmap viewer is provided at https://jades.idies.jhu.edu, with a version tailored for the public at https://jades.idies.jhu.edu/public.
-Observations of the Cycle 2 program 3215 are underway at the time of this writing (Oct.
-2023) and will be immediately public.
 The power of JWST to reveal the early history of galaxies with infrared imaging and spectroscopy is both transformative and inspiring.
 Many compelling observing programs have been performed or are underway, each with their own nuances and opportunities.
-We hope that the extraordinary depth and filter complement of the JADES Origins Field will provide a superb new view of galaxies in their youth.
+We hope that the extraordinary depth and filter complement of the JADES Origins Field provides a superb new view of galaxies in their youth.